### v1

# JUHAN ARU, PHILÉMON BORDEREAU

Abstract.
One way to uniquely define Schramm-Loewner Evolution (SLE) in multiply connected domains is to use the restriction property. This gives an implicit definition of a $\sigma$-finite measure on curves; yet it is in general not clear how to construct such measures nor whether the mass of these measures, called the partition function, is finite.
We provide an explicit construction of the such conformal restriction SLEs in multiply connected domains when $\kappa=4$ using the Gaussian Free Field (GFF). In particular, both when the target points of the curve are on the same or on distinct boundary components, we show that there is a mixture of laws of level lines of GFFs that satisfies the restriction property. This allows us to give an expression for the partition function of $\mathrm{SLE}_{4}$ on multiply connected domains and shows that the partition function is finite, answering the question raised in [Law09, Law11].
In a second part, we provide a second construction of $\mathrm{SLE}_{\kappa}$ in multiply-connected domains for the whole range $\kappa\in(8/3,4]$; specific, however, to the case of the two target points belonging to the same boundary components. This is inspired by [WW13] and consists of a mixture of laws on curves obtained by following $\operatorname{CLE}_{\kappa}$ loops and restriction hulls attached to parts of the boundary of the domain. In this case as well, we obtain as a corollary the finiteness of the partition function for this type of $\mathrm{SLE}_{\kappa}$.
Introduction
Schramm-Loewner Evolution $\mathrm{SLE}_{\kappa}$, $\kappa\geq 0$ is a one-dimensional family of conformally invariant probability measures on non-crossing curves in domains of the complex plane [Sch00] that are either proved or conjectured to describe the scaling limit of interfaces of a variety of planar statistical physics models, e.g. [Smi10, LSW04, SS09, DCLQ21, CDCH+14, KW10, She06]. Another active area of research since the introduction of SLE has been to identify and describe its connections with Conformal Field Theory (CFT), where $\mathrm{SLE}_{\kappa}$ is expected to be connected with CFTs of central charge $\mathbf{c}=\mathbf{c}(\kappa)=(6-\kappa)(3\kappa-8)/(2\kappa)$, see e.g. [Pel19] for a review.
Schramm constructed (chordal) $\mathrm{SLE}_{\kappa}$ in simply connected domains as a solution to Loewner’s equation driven by a multiple of Brownian motion $\sqrt{\kappa}B$ and showed via a short argument that the resulting law is the unique law on simple curves (modulo increasing time-reparametrization) that satisfies conformal invariance and a domain Markov property [Sch00]. Neither this construction nor the characterisation generalise directly to multiply connected domains as the conformal class changes with the growth of the curve and therefore the remaining slit domain cannot be related to the initial domain via conformal transformations–in other words, due to the non-triviality of the moduli space for more general domains, in order to identify a family of laws on curves indexed by the moduli space, one needs to specify how these laws between different conformal classes compare.
Still, variants of $\mathrm{SLE}_{\kappa}$ have been constructed in doubly connected domains via the annulus Loewner equation using the modulus of the slit domain as the time parametrization [Zha04, Zha08, HBB10, HLD08], and in higher connectivity using further generalizations of the Loewner equation [BF08, BF04]. In these latter works one also needs to keep track of the evolution of the moduli as the curve grows to obtain a Markov property and this has a downside that the natural conditions do not allow to single out an unique law on the curves, but rather one identifies a whole family of possible curves as $\mathrm{SLE}_{\kappa}$. A different approach that does single out an unique law of $\mathrm{SLE}_{\kappa}$ was suggested by Lawler [Law09] and is described next.
Lawler’s definition of $\mathrm{SLE}$ in multiply connected domains via conformal restriction
In [LSW03], it was noted that for $\kappa\leq 4$ (the values of $\kappa$ for which the resulting curve is a.s. simple [RS05]), the law $\nu_{\mathbb{H}\setminus K,\kappa}(0,+\infty)$ of $\mathrm{SLE}_{\kappa}$ from 00 to $\infty$ with $K$ compact, $\mathrm{dist}(K,0)>0$ and $\mathbb{H}\setminus K$ connected is absolutely continuous with respect to $\nu_{\mathbb{H},\kappa}(0,+\infty)$ and that their Radon-Nikodym derivative is explicitly given by
(1) $\frac{\mathrm{d}\nu_{\mathbb{H}\setminus K,\kappa}(0,+\infty)}{\mathrm{d}\nu_{\mathbb{H},\kappa}(0,+\infty)}(\eta)=\mathbf{1}_{\eta\subset\mathbb{H}\setminus K}\varphi^{\prime}(0)^{-h}\exp(\frac{\mathbf{c}}{2}m_{\mathbb{H}}(\eta,K)),$
where $\varphi:\mathbb{H}\setminus K\to\mathbb{H}$ is a conformal equivalence chosen such that $\varphi(0)=0,\varphi(z)\sim z$ as $z\to\infty$, and $h=h_{\kappa}=(6-\kappa)/(2\kappa)$ is a constant that can be related to the weight of a boundary condition changing operator in a CFT [Car05, Pel19, Dub15]. The term $m_{\mathbb{H}}(\eta,K)$ corresponds to the Brownian measure (see [LW04] for a definition thereof) on loops in $\mathbb{H}$ that intersect $K$ and $\eta$. This restriction property is one motivation for seeing the laws $\nu_{D,\kappa}(x,y)$ no longer as probability measures, but rather as measures of total mass [Law09, Law11]
$Z_{D,\kappa}(x,y)=\begin{cases}H_{\partial D}(x,y)^{h}&\text{if }x\neq y\in D\setminus\left\{\infty\right\}\\ \absolutevalue{\varphi^{\prime}(x)}^{h}&\text{if }y=\infty\end{cases}$
where $H_{\partial D}$ denotes the boundary Poisson kernel (when $\partial D$ is sufficiently regular at $x$ and $y$) and $\varphi:D\to\mathbb{H}$ is a conformal equivalence with $\varphi(x)=0$, $\varphi(z)\sim z$ as $z\to\infty$. This total mass is called the partition function of the chordal $\mathrm{SLE}_{\kappa}$ and it is defined up to a constant such that $Z_{\mathbb{H}}(0,+\infty)=1$. The measures $\nu$ then satisfy a form of conformal covariance
$\varphi_{\ast}\nu_{D,\kappa}(x,y)=\absolutevalue{\varphi^{\prime}(x)}^{h}\absolutevalue{\varphi^{\prime}(y)}^{h}\nu_{\varphi(D),\kappa}(\varphi(x),\varphi(y))$
for $x,y\in\partial D$ and $\varphi:D\to\varphi(D)$ a conformal equivalence (smooth at $x$ and $y$), suggesting that these partition functions may be also interpreted as operators in a Conformal field theory [Dub15] and shows they satisfy BPZ equations. This raises a natural question: are the partition functions $Z_{D,\kappa}(x,y)$ positive and finite? If this were the case, the measures could be normalized and the resulting probability measures would be conformally invariant.
Furthermore, Lawler suggested that the restriction property (1) could be used to implicitly but uniquely define $\mathrm{SLE}$ on multiply connected domains $D$ as follows (see Definition 3 for more precision):
Definition 3 (Conformal restriction $\mathrm{SLE}_{\kappa}$ in multiply-connected domains).
For $0\leq\kappa\leq 4$, chordal $\mathrm{SLE}_{\kappa}$ is the family of laws $\left\{\nu_{D,\kappa}(x,y)\right\}_{(D,x,y)}$, where $D\subset\mathbb{C}$ is a finitely-connected domain with smooth boundary at $x$ and $y$, such that
  • the renormalized probability measures are conformally invariant, and
  • the restriction of $\nu_{D,\kappa}(x,y)$ to curves remaining in a simply connected domain $D^{\prime}\subset D$ is absolutely continuous with the $\mathrm{SLE}_{\kappa}$ measure $\nu_{D^{\prime},\kappa}(x,y)$ in $D^{\prime}$, with Radon-Nikodym derivative
    (2) $\frac{\mathrm{d}\nu_{D,\kappa}(x,y)|_{D^{\prime}}}{\mathrm{d}\nu_{D^{\prime},\kappa}(x,y)}(\eta)=\mathbf{1}_{\eta\subset D^{\prime}}\exp(-\frac{\mathbf{c}}{2}m_{D}(\eta,D^{\prime}\setminus D)).$
Because these Radon-Nikodym derivatives are consistent, one can see (for instance via a Carathéodory extension argument) that Equation (2) indeed implicitly defines a ($\sigma$-finite) measure on curves on any finitely connected domain of the complex plane, since the additivity structure of the Brownian loop measure makes the restriction to different domains compatible [Law11]. Moreover, these measures are unique up to a multiplicative constant that one can fix by requiring that $Z_{\mathbb{H}}(0,+\infty)=1$, and satisfy the domain Markov property that is expected of SLE, as well as a form of reversibility. It is further shown in [Law11] that $Z_{D,\kappa}(x,y)$ is positive, and finite either when $\kappa\leq 8/3$ (corresponding to the regime $\mathbf{c}\leq 0$) or when $\kappa\leq 4$ and $D$ is one-connected. The main aim of this paper is to give explicit constructions of the laws satisfying the definition above, and to use those to prove that the partition function is non-trivial.
Before describing our results, let us recall two other ways of constructing $\mathrm{SLE}$ in simply-connected domains, that we will generalize. There is yet a third fruitful way via Liouville quantum gravity and welding of random surfaces [SY23] that we will not treat in this current paper and that is a priori harder to use for the case $\mathrm{SLE}_{4}$ that is the main topic here.
Other constructions of $\mathrm{SLE}$ in simply connected domains
In the case of simply connected domains, $\mathrm{SLE}$ curves can be seen to appear from other conformally invariant objects:
  • The Gaussian Free Field (GFF) corresponds to a centered Gaussian process with covariance operator formally given by Green’s function. Its appearance as a scaling limit for distributions suggests a connection with $\mathrm{SLE}$, which was first discovered in the seminal paper [SS09]. While $\mathrm{SLE}_{\kappa}$ for $\kappa<4$ can be understood as flow lines of the GFF [MS16, She16], it is shown that $\mathrm{SLE}_{4}$ arises as a level-line of the GFF: both as a scaling limit of a level-line of the discrete GFF in [SS09], but also as a purely continuum interface in [SS13]. More precisely, a curve $\eta\sim\nu_{D,4}(x,y)$ in a simply-connected domain $D$ with $x,y\in\partial D$ can be coupled with a Gaussian Free Field $\Phi_{D}$ with mean corresponding to the harmonic extension of the boundary values $-\lambda:=-\sqrt{\pi/8}$ (resp. $+\lambda$) on the counterclockwise arc $(yx)$ (resp. $(xy)$), so that conditionally on $\eta$, $\Phi_{D}$ is equal in law to a sum of two independent GFFs on the left (resp. on the right) of $D\setminus\eta$, with constant mean equal to $-\lambda$ (resp. $+\lambda$). Further works explore variations of this results for other types of SLEs [PW17, HBB10, IK13], and prove the existence of level-lines for GFFs in multiply connected domains [ASW19].
  • The (one-sided) restriction measures and the Conformal Loop Ensembles (CLE) are one-parameter families of measures on hulls (resp. loops) satisfying a form of conformal Markov property. Both can be constructed from Brownian motion: for $D\subset\mathbb{C}$ a bounded domain with $x,y\in\partial D$, the filling (relatively to $(yx)$) of a Poisson point process of Brownian excursions on the boundary arc $(yx)\subset\partial D$ of intensity $\alpha$ yields a one-sided restriction measure of parameter $h=h(\alpha)$ (see Section 3.1 and [LSW03]), whereas the collection of outer boundaries of clusters of loops generated by a Poisson point process of intensity $c$ with respect to the Brownian loop measure corresponds to a $\operatorname{CLE}_{\kappa}$ with $\kappa=\kappa(c)$ (see Section 3.1 and [SW12]). In [WW13], it was shown that for $D\subset\mathbb{C}$ simply connected with $x\neq y\in\partial D$ and $\kappa\in(8/3,4]$, if one samples independently a restriction measure on $(xy)$ of parameter $h_{\kappa}=(6-\kappa)/(2\kappa)$ and a $\operatorname{CLE}$ of parameter $\kappa$, and then proceeds to follow the outer boundary of the restriction sample along with the $\operatorname{CLE}$ loops it intersects, then the curve obtained is distributed like $\mathrm{SLE}_{\kappa}$ in $D$.
Summary of results
In this article, we extend these constructions to the case of multiply connected domains in a way that allows us the study the conformal restriction $\mathrm{SLE}_{\kappa}$ as defined in Definition 3.
In a first part, we provide a construction specific to $\mathrm{SLE}_{4}$ that relies on the connection with the GFF, both when the two target points belong to the same boundary component (non-crossing case) and different boundary components (crossing-case). Loosely speaking, the results for $\mathrm{SLE}_{4}$ can be stated as follows:
  • $\star$
    The law of conformal restriction $\mathrm{SLE}_{4}$ in any multiply connected domain can be obtained from a mixture of level-lines of the GFF via appropriate topological conditionings. (See Theorem 913)
  • $\star$
    The partition function of conformal restriction $\mathrm{SLE}_{4}$ can be identified, and is finite and positive. (See Corollary 1215)
Somewhat surprisingly, the conformal restriction $\mathrm{SLE}_{4}$ cannot be seen as a level line of a single GFF (see Remark 2.2), except in the crossing-case in the annulus, where it can be interpreted as a level line of a compactified GFF. A bit more precisely, the mixture of laws can be understood as conditioning the final curve on the boundary components it leaves to the left, then sampling a level-line for a GFF whose mean is $-\lambda$ on those boundary components, $+\lambda$ on the others, and conditioning on the level-line to indeed leave those boundary components to the left.
Except for topological discrepancies, the proofs for both cases then follow the same strategy: the law of the level-line $\eta$ of a GFF $\Phi_{D}$ in a domain $D$ with suitable boundary conditions conditioned to remain in a smaller domain $D^{\prime}\subset D$ can be seen as a level-line for a GFF in $D^{\prime}$ with random boundary conditions given by the values $\xi$ of $\Phi_{D}$ on $\partial D^{\prime}$. We first compare this law to $\nu_{D^{\prime},4}(x,y)$ for $\xi$ fixed in Propositions 11 and 17. This builds on the techniques developed in [ALS20b, ALS22] to compare the laws of local sets of GFFs under changes of boundary conditions: Proposition 11 can be seen as a variant of Proposition 3.3 in [ALS22] for more general boundary conditions, but specific to SLEs. In a second step, we compute the average of this Radon-Nikodym derivative over $\xi$, conditionally on $\eta\subset D^{\prime}$. This is a Gaussian integration, and its result consists of a determinant term corresponding to a mass of loops by Proposition 2.2 in [Dub09], and (an exponential of) a Dirichlet energy which we identify as the partition function of the SLE. Equivalently, it can be seen as the inverse extremal distance between parts of the boundary of the domain.
In a second part, we provide another construction of $\mathrm{SLE}_{\kappa}$, for the whole range $\kappa\in(8/3,4]$, but specific to the non-crossing case:
  • $\star$
    In any multiply-connected domain and for any $\kappa\in(8/3,4]$ we can construct the conformal restriction $\mathrm{SLE}_{\kappa}$ with the endpoints on different boundary components as a mixture of laws obtained from Brownian excursions and CLE via topological conditioning. (See Theorem 19)
  • $\star$
    In the setup just described, one can give an expression for the partition function and show that it is finite. (See Corollary 20)
More precisely, conditionally on the boundary components on the left of the curve, one samples a restriction measure on these boundary components plus the arc $(yx)$ and an independent $\operatorname{CLE}$, and follows the right boundary together with the $\operatorname{CLE}$ loops encountered along the way. To verify that the restriction property (2) is satisfied, it is sufficient to compute the Radon-Nikodym derivative of this law against that in a domain $D\setminus K$, $K$ a hull at positive distance from $(yx)$: the contribution from the $\operatorname{CLE}$ provides the term $\exp(\mathbf{c}\cdot m_{D}(\eta,K)/2)$, whereas the one for the restriction measures yields a normalizing constant that we identify as the partition function of the $\mathrm{SLE}$ (in the simply connected case, this term would be $H_{\partial D}(x,y)^{h}$). We were however not able to extend this construction to the crossing case, which appears to be more delicate - it is not clear how to use the geometric construction via Brownian excursions and CLE to define a crossing $\mathrm{SLE}_{\kappa}$, even in the simplest case of $\mathrm{SLE}_{4}$.
The article is structured as follows. Section 1 reviews standard background on complex analysis tools and $\mathrm{SLE}$. We move the reminders about the GFF and the restriction measures and $\operatorname{CLE}$ to the second and third sections where these tools appear specifically. In Section 2, we construct $\mathrm{SLE}_{4}$ first in the non-crossing case then in the crossing case, from the GFF. Section 3 provides another construction of the $\mathrm{SLE}_{\kappa}$ measure for $\kappa\in(8/3,4]$ via the restriction measures and $\operatorname{CLE}$, but specifically to the non-crossing setting.
Acknowledgements
Both authors are supported by Eccellenza grant 194648 of the Swiss National Science Foundation and are members of NCCR Swissmap.
1. Preliminaries
1.1. Notations
We establish notation and definitions that will be used throughout this text:
  • $\mathbb{D}$ denotes the unit disk in $\mathbb{C}$, $\mathbb{H}$ the upper half plane, and for $p>0$, $\mathbb{A}_{p}$ denotes the annulus $\mathbb{A}_{p}:=\left\{z\in\mathbb{C}:e^{-p}<\absolutevalue{z}<1\right\}$. A domain is a strict open subset of $\mathbb{C}$. We denote $D\subset\mathbb{C}$ for this strict inclusion.
  • Given $x\neq y\in\partial\mathbb{D}$, a domain $D\subset\mathbb{D}$ is an (non-crossing) admissible domain if it is finitely connected (in particular, connected) with the connected components of $\partial D$ given by $\mathcal{B}\cup\bigcup_{i=1}^{n}\mathcal{B}_{i}$, $n\geq 0$, with $\mathcal{B}$ and each $\mathcal{B}_{i}$ given by a Jordan curve. We further ask that $x,y\in\mathcal{B}\subset\partial D$ and that $D$ agrees with $\mathbb{D}$ in a neighborhood of $x$ and $y$. We denote by $(xy)$ or $\mathcal{B}_{r}$ (resp. $(yx)$ or $\mathcal{B}_{l}$) the counterclockwise arc from $x$ to $y$ (resp. from $y$ to $x$) in $\mathcal{B}$ (see also Figure 1).
    On the other hand, given $x\in\partial\mathbb{D}$ and $y\in\mathbb{D}$, a (crossing) admissible domain is a finitely connected domain $D\subset\mathbb{D}$ with $\partial D=\mathcal{B}_{\text{in}}\cup\mathcal{B}_{\text{out}}\cup\bigcup_{i=1}^{n}\mathcal{B}_{i}$, $n\geq 0$, with each component of $\partial D$ given by a Jordan curve. We also require that $x\in\mathcal{B}_{\text{out}}$ with $\mathcal{B}_{\text{out}}$ and $\partial\mathbb{D}$ agreeing locally around $x$, and that $y\in\mathcal{B}_{\text{in}}$ is analytic, i.e. that $\partial D$ is smooth at $y$.
    The requirement that boundary components correspond to Jordan curves is to avoid punctured domains. Furthermore, note that every admissible domain is conformally equivalent to a circle domain of the unit disk, i.e. such that $\mathcal{B}=\partial\mathbb{D}$, with each $\mathcal{B}_{i}$ corresponding to a circle in $\partial\mathbb{D}$: in particular we may assume that all boundary components are smooth, because the $\mathrm{SLE}$ measures that we construct can be transported by conformal equivalences.
  • Similarly, a domain $D^{\prime}$ is called a test domain (relatively to $(D,x,y)$ an admissible domain, crossing or non-crossing) if $D^{\prime}\subset D$, $D^{\prime}$ is simply-connected with $x,y\in\partial D^{\prime}=:\mathcal{B}^{\prime}$ and $D^{\prime}$ agrees with $D$ in a neighborhood of $x$ and $y$. We denote by $\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}^{\prime}$) the part of the counterclockwise arc from $y$ to $x$ (resp. from $x$ to $y$) in $\mathcal{B}^{\prime}$ that lies in $D$.
  • For $\eta:[0,t_{\eta}]\to D$ a curve in a domain $D$ with $0\leq t_{\eta}\leq+\infty$, we write $\eta_{t}:=\eta([0,t])$ for $t\in[0,t_{\eta}]$. We abuse notation and write $\eta\subset D$ if $\eta(0,t_{\eta})\subset D$, allowing for $\eta(0),\eta(t_{0})\in\partial D$, which we call a curve from $\eta(0)$ to $\eta(t_{\eta})$ in $D$ (notice that the direction of time in the parametrization of the curve is implicit here). If $t<t_{\eta}$, we write $\eta_{t}\subset D$ when $\eta(0,t]\subset D$. If $\eta(0)$ and $\eta(t_{\eta})$ belong to the same boundary component, it separates $D$ in two parts we call $D_{l}$ or $(D\setminus\eta)_{\text{left}}$ (resp. $D_{r}$ or $(D\setminus\eta)_{\text{right}}$) the left (resp. right) connected component of $D\setminus\eta$ relatively to the orientation of $\eta$.
  • For $D$ a simply-connected domain and $x\neq y\in\partial D$, we denote by $g_{D}$ the harmonic extension of the boundary values $-\lambda:=-\sqrt{\pi/8}$ (resp- $+\lambda$) on $(yx)$ (resp. on $(xy)$).
  • For $D$ a domain, we let $(\cdot,\cdot)_{\nabla(D)}$ denote the Dirichlet inner product
    $(f,g)_{\nabla(D)}=\int_{D}\nabla f\cdot\nabla g$
    for $f,g$ differentiable. For $\delta\subset D$ a smooth path or curve, we denote by $(\cdot,\cdot)_{\delta}$ the scalar product on $L^{2}(\delta)$. We will often abuse notation and write $(f,g)_{\delta}$ when $f$ is a distribution on $\delta$ tested against a smooth enough function $g$.
1.2. Poisson kernels and operators
We review some definitions and identities between Green’s function and Poisson kernels.
  • The Green’s function in a domain $D$ at $z\neq w\in D$ is the density at $w$ of a standard Brownian motion started at $z$, killed upon exiting $D$. This is normalized so that $G_{\mathbb{D}}(0,z)=-(2\pi)^{-1}\log\absolutevalue{z}$ |. We stress that this is a choice of normalization that might differ to some papers in the field, where the Green’s function would be defined as $\tilde{G}_{D}(0,z)=-\log\absolutevalue{z}$ |. It is convenient for us here as it yields $G_{D}=(-\Delta)^{-1}$ distributionally, without extra multiplicative terms. Note that the choice for the multiplicative constant of $G_{D}$ affects the value of $\lambda$ (choosing $G_{D}(x,y)=-c\log\absolutevalue{x-y}+o(1)$ as |$\absolutevalue{x-y}\to 0$ corresponds to $\lambda=\pi\sqrt{c}/2$).
  • For $D$ a domain, a point $w\in\partial D$ is $\partial D$-analytic if there exists a neighborhood $N$ in $\mathbb{C}$ and $\varphi:N\to\mathbb{C}$ conformal with $\varphi(w)=0,\varphi(N\cap D)=\varphi(N)\cap\mathbb{H}$. In an admissible domain $(D,x,y)$ as above, $x$ and $y$ are $\partial D$-analytic. For $z\in D$ and $w$ $\partial D$-analytic, we let $H_{D}(z,w)$ denote the Poisson kernel, defined as the density at $w$ of the measure induced by the exit point of $D$ of a Brownian motion started at $z$, i.e. for all $V\in\partial D\cap N$,
    $\mathbb{P}\left[B_{T_{D}}^{z}\in V\right]=\int_{V}H_{D}(z,w^{\prime})\mathrm{d}w^{\prime},$
    well-defined as $w$ is $\partial D$-analytic. Under this normalization,
    $H_{\mathbb{H}}(x+iy,0)=\frac{1}{\pi}\frac{y}{x^{2}+y^{2}}.$
  • For $z,w$ both $\partial D$-analytic, we let $H_{\partial D}(x,y)$ denote the boundary or excursion Poisson kernel, defined by
    $H_{\partial D}(z,w)=\partial_{\nu_{z}}H_{D}(z,w)=\partial_{\nu_{w}}H_{D}(z,w),$
    with $\nu_{z}$ the inward normal vector at $z$. We stress again that the choice of normalization for the Poisson kernel carries over to the boundary Poisson kernel.
These three functions satisfy the following conformal covariance property: for $\varphi:D\to\varphi(D)$ a conformal equivalence,
$G_{D}(x,y)$ $=G_{\varphi(D)}(\varphi(x),\varphi(y)),$ $x,y\in D$
$H_{D}(x,y)$ =|$=\absolutevalue{\varphi^{\prime}(y)}H_{\varphi(D)}(\varphi(x),\varphi(y)),$ $x\in D,y\in\partial D$
$H_{\partial D}(x,y)$ =|$=\absolutevalue{\varphi^{\prime}(x)}\absolutevalue{\varphi^{\prime}(y)}H_{\partial\varphi(D)}(\varphi(x),\varphi(y)),$ $x,y\in\partial D$
with $x$ and $\varphi(x)$ analytic on $\partial D$ and $\partial\varphi(D)$ whenever $x\in\partial D$, same for $y$, so that the quantities are well-defined.
We now relate the Green’s function and the Poisson kernels to the mass of Brownian measures [LW04, Law08]. We can decompose the law $\mu(z,t)$ of a Brownian motion $B$ running at speed $2$ (this is to later match the normalization of the Green’s function) started at $z$, of time-duration $t$ as
$\mu(z,t)=\int_{\mathbb{C}}\mu(z,w,t)\mathrm{d}w,$
and identify $\mu(z,w,t)$ as $(4\pi t)^{-1}\exp(-\absolutevalue{z-w}^{2}/(4t))$ times the probability measure corresponding to a Brownian bridge of time-duration $t$ from $z$ to $w$. In a domain $D\subset\mathbb{C}$, the Brownian path measure is defined by restriction to paths that remain in $D$: as such, the Brownian measures verify a restriction property, and a form of conformal covariance that is inherited from the conformal invariance of Brownian motion. Green’s function then naturally appears as the mass of the time-integrated measure
$\mu_{D}(z,w)=\int_{0}^{+\infty}\mu_{D}(z,w,t)\mathrm{d}t.$
The measures $\mu_{D}(z,z)$ are infinite for $z\in D\subset\mathbb{C}$, with divergences coming from small loops (this corresponds to the fact that $G_{D}(z,w)\sim(2\pi)^{-1}\log\absolutevalue{z-w}$ | as $z\to w$). Furthermore, they are measures on rooted loops, and forgetting the root leads to the Brownian loop measures
$\mu_{D}^{\text{loop}}=\int_{D}\int_{0}^{+\infty}\frac{1}{2t^{2}}\mu_{D}^{\#}(z,z,t)\mathrm{d}t\mathrm{d}z,$
with $\mu_{D}^{\#}(z,z,t)$ corresponding to the law of a Brownian bridge from $z$ to $z$ of time duration $t$ on the event that it stays in $D$. The importance of these measures comes from the fact that they satisfy the restriction property, but also–less obviously–conformal invariance [LW04]. For $K_{1},K_{2}$ two compacts of $D$ such that $\mathrm{dist}(K_{1},K_{2})>0$, the mass
$m_{D}(K_{1},K_{2})=\mu_{D}^{\text{loop}}(\left\{\omega\text{ loop in }D:\omega\cap K_{1}\neq\emptyset,\omega\cap K_{2}\neq\emptyset\right\})$
of loops in $D$ that intersect both $K_{1}$ and $K_{2}$ is finite and also a conformal invariant: for $\varphi:D\to\varphi(D)$ a conformal equivalence, $m_{D}(K_{1},K_{2})=m_{\varphi(D)}(\varphi(K_{1}),\varphi(K_{2}))$. This quantity is also related to the (zeta-regularized) determinant of the Laplacian, see Proposition 2.1 in [Dub09].
The Poisson and boundary Poisson kernels also describe the masses of the interior-to-boundary and boundary-to-boundary (excursions) Brownian measures. If $x,y\in\partial D$ are $\partial D$-analytic, then one can define the boundary-to-boundary or excursion measure as
$\mu_{\partial D}(x,y):=\lim_{\varepsilon\downarrow 0}\frac{1}{2\varepsilon^{2}}\mu_{D}(x+\varepsilon n_{x},y+\varepsilon n_{y}),$
a measure on paths from $x$ to $y$ in $D$. It can be realized as a measure of mass $H_{\partial D}(x,y)$ describing a Brownian excursion from $x$ to $y$, see once again [LSW03]. If $D^{\prime}\subset D$ agrees with $D$ around $x$ and $y$, then the definition of $\mu_{\partial D}(x,y)$ as a limit of interior Brownian measures readily shows that $\mu_{\partial D}(x,y)$ restricted to paths $\eta\subset D^{\prime}$ coincides with $\mu_{\partial D^{\prime}}(x,y)$.
We can also see the Poisson kernel and Green’s function as differential operators. For instance, for $\mathcal{B}_{1},\mathcal{B}_{2}$ disjoint smooth curves in a domain $D$, define $H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}:L^{2}(\mathcal{B}_{1})\to L^{2}(\mathcal{B}_{2})$ as
$H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}f=\int_{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}(\cdot,x)f(x)\mathrm{d}x.$
$(G_{D})_{\mathcal{B}_{i}}^{\mathcal{B}_{j}}$ can also be seen as an operator in the same way, but in this case allowing for $i=j$:
$((G_{D})_{\mathcal{B}_{i}}^{\mathcal{B}_{j}}f)(y)=\int_{\mathcal{B}_{i}}G_{D}(y,x)f(x)\mathrm{d}x\quad\text{ for }f\in L^{2}(\mathcal{B}_{i}),y\in\mathcal{B}_{j}.$
Similarly, we define $(H_{\partial D}-H_{\partial D\setminus\mathcal{B}_{2}})^{\mathcal{B}_{1}}:L^{2}(\mathcal{B}_{1})\to L^{2}(\mathcal{B}_{1})$ as
$(H_{\partial D}-H_{\partial D\setminus\mathcal{B}_{2}})^{\mathcal{B}_{1}}f=\int_{\mathcal{B}_{1}}(H_{\partial D}-H_{\partial D\setminus\mathcal{B}_{2}})(\cdot,x)f(x)\mathrm{d}x.$
We stress that only the difference of the boundary Poisson kernels is integrable (as for instance $H_{D}(x,y)\asymp(x-y)^{-2}$). We will often drop the superscripts when the context makes it clear.
We first gather two useful identities.
Lemma 1 (Some identities for loop measures).
Let $\mathcal{B}_{1},\mathcal{B}_{2}$ be disjoint smooth simple curves in $D$ with $\mathcal{B}_{1}$ smooth, and $x\neq y\in\mathcal{B}_{1}$. It holds that
(3) $(G_{D}-G_{D\setminus\mathcal{B}_{2}})(x,y)=\int_{\mathcal{B}_{2}}H_{D\setminus\mathcal{B}_{2}}(x,\mathrm{d}z)\int_{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}(z,\mathrm{d}w)G_{D}(w,y),$
(4) $\int_{\mathcal{B}_{2}}H_{D\setminus\mathcal{B}_{2}}(x,\mathrm{d}z)H_{D\setminus\mathcal{B}_{1}}(z,x)=\int_{\mathcal{B}_{1}}(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})(x,\mathrm{d}z)G_{D\setminus\mathcal{B}_{2}}(z,x).$
We might sometimes write more succintly these equations as
$(G_{D}-G_{D\setminus\mathcal{B}_{2}})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}$ $=H_{D\setminus\mathcal{B}_{2}}^{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}(G_{D})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}},$
$H_{D\setminus\mathcal{B}_{2}}^{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}$ $=(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})^{\mathcal{B}_{1}}(G_{D})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}.$
Proof.
The left-hand side in the first equation corresponds to the mass of Brownian paths from $x$ to $y$ in $D$ that hit $\mathcal{B}_{2}$. One can decompose such a path at its first hitting time of $\mathcal{B}_{2}$, and at its following first hitting time of $\mathcal{B}_{1}$: the mass of the measure disintegrated in this way is exactly the right-hand side.
The second equation follows similarly by the Markov property and reversibility of the Brownian measure: the left-hand side counts Brownian loops attached at $\mathcal{B}_{1}$ that ‘enter’ immediately $D\setminus\mathcal{B}_{1}$, hit $\mathcal{B}_{2}$, and come back to their starting point in $D$, which can be decomposed as an excursion from $\mathcal{B}_{1}$ to $\mathcal{B}_{1}$ in $D\setminus\mathcal{B}_{1}$ hitting $\mathcal{B}_{2}$ and a path to the starting point in $D\setminus\mathcal{B}_{2}$. ∎
Finally, we need a Proposition from [Dub09] that will be relevant in simplifying computations of Gaussian integrations:
Proposition 2 (Proposition 2.2 in [Dub09]).
In the setting above, the following identity holds:
$\exp(-m_{D}(K_{1},K_{2}))=\mathrm{det}_{F}(I_{\mathcal{B}_{1}}-H_{D\setminus\mathcal{B}_{2}}^{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}),$
where $\mathrm{det}_{F}$ stands for the Fredholm determinant.
$I_{\mathcal{B}_{1}}$ stands here for the identity operator on $L^{2}(\mathcal{B}_{1})$. We will leave the notation $1_{\mathcal{B}_{1}}$ for the constant function in $L^{2}(\mathcal{B}_{1})$ equal to one.
1.3. Dirichlet energies
In what follows, we will often be brought to consider diverging Dirichlet energies of harmonic functions that are locally constant on the boundary of a domain. Consider $D$ a regular bounded simply connected domain again with $x\neq y\in\partial D$, with boundary $\mathcal{B}_{l}\cup\mathcal{B}_{r}=(yx)\cup(xy)$ as in the notation section. Let $u_{D}$ denote the harmonic extension of the boundary values $-1$ on $\mathcal{B}_{l}$, $+1$ on $\mathcal{B}_{r}$. The Dirichlet energy $\norm{u_{D}}_{\nabla(D)}:=\norm{\nabla u_{D}}_{D}$ is clearly ill-defined, but we can regularize it at the discontinuity points of $u_{D}$:
$\norm{u_{D}}_{\nabla(D),\text{reg}}^{2}:=\lim_{\varepsilon_{1}\to 0,\varepsilon_{2}\to 0}\int_{z\in D,\absolutevalue{z-x}>\varepsilon_{1},\absolutevalue{z-y}>\varepsilon_{2}}\absolutevalue{\nabla u_{D}}^{2}\mathrm{d}z+\frac{4}{\pi}(\log\varepsilon_{1}+\log\varepsilon_{2}).$
It is easy to see that the limit exists. Nonetheless, in many relevant situations below, we will consider differences of such diverging Dirichlet energies to which it is possible to provide a meaning without renormalization. Consider in the above set-up $D^{\prime}\subset D$ another test domain such that $\mathrm{dist}(\mathcal{B}_{l},D\setminus D^{\prime})>0$. One might formally write
$\int_{D}\absolutevalue{\nabla u_{D}}^{2}$ $=-\int_{\mathcal{B}_{l}}\partial_{\nu}u_{D},$
$\int_{D^{\prime}}\absolutevalue{\nabla u_{D^{\prime}}}^{2}$ $=-\int_{\mathcal{B}_{l}}\partial_{\nu}u_{D^{\prime}},$
and define
$\norm{u_{D^{\prime}}}_{\nabla(D^{\prime})}-\norm{u_{D}}_{\nabla(D)}:=\int_{\mathcal{B}_{l}}\partial_{\nu}(u_{D}-u_{D^{\prime}})$
where the normal derivative is well-defined and the integral is finite since $u_{D}-u_{D^{\prime}}$ is harmonic in $D^{\prime}$ with boundary values 00 on $\partial D\cap\partial D^{\prime}$ and $u_{D}-1\geq 0$ on $\delta$. One can also replace $u_{D},u_{D^{\prime}}$ by their expression involving the Poisson kernel:
$u_{D}(\cdot)=1-\int_{\mathcal{B}_{l}}H_{D}(\cdot,\mathrm{d}y),\quad u_{D^{\prime}}(\cdot)=1-\int_{\mathcal{B}_{l}}H_{D^{\prime}}(\cdot,\mathrm{d}y)$
so that the definition above becomes (using that $\partial_{\nu_{z}}H_{D}(z,w)=-H_{\partial D}(z,w)$ here, with $\nu_{z}$ the outwards normal vector)
$\norm{u_{D^{\prime}}}_{\nabla(D^{\prime})}-\norm{u_{D}}_{\nabla(D)}$ $:=\iint_{\mathcal{B}_{l}\times\mathcal{B}_{l}}(H_{\partial D}-H_{\partial D^{\prime}})(\mathrm{d}z,\mathrm{d}w)$
$=(1_{\mathcal{B}_{l}},(H_{\partial D}-H_{\partial D^{\prime}})1_{\mathcal{B}_{l}})_{\mathcal{B}_{l}}.$
This also corresponds to the mass of Brownian excursions from $\mathcal{B}_{l}$ that hit $D\setminus D^{\prime}$. In particular, the resulting expression is conformally invariant. The difference of these Dirichlet energies can also be seen as the difference of the inverse extremal distance $M(\mathcal{B}_{l},\mathcal{B}_{r}^{\prime})-M(\mathcal{B}_{l},\mathcal{B}_{r})$, see [Ahl10, Theorem 4-5]. The formula above also clearly generalizes when $D$ is a general admissible domain (possibly multiply connected), and $\partial D=\mathcal{B}_{-}\cup\mathcal{B}_{+}$ with $\mathcal{B}_{-}$, $\mathcal{B}_{+}$ both containing an arc from $x$ to $y$, as well as the other boundary components of $\partial D$.
1.4. $\mathrm{SLE}$ measures in simply connected domains
Let us recall the construction of $\mathrm{SLE}_{\kappa}$ in the upper half-plane, before reviewing them as non-probability measures. Given $\kappa>0$, let $B_{\kappa\cdot}$ be a one-dimensional Brownian motion running at speed $\kappa$, and $(g_{t}(z))_{t\geq 0}$ be the solution to the following chordal Loewner equation:
$\partial_{t}g_{t}(z)=\frac{2}{g_{t}(z)-B_{\kappa t}},\quad g_{0}(z)=z,\quad\quad z\in\mathbb{H}.$
For each $z\in\mathbb{H}$, the flow above is well-defined until $T(z)=\inf\left\{t\geq 0:g_{t}(z)-B_{\kappa t}=0\right\}$. The resulting hulls $K_{t}:=\left\{z\in\mathbb{H}:T(z)\leq t\right\}$ almost surely correspond to simple curves for $0\leq\kappa\leq 4$, and are generated by non-crossing curves for $4<\kappa$ (i.e. there exists $\gamma$ such that $K_{t}$ is equal to the unbounded connected component of $\mathbb{H}\setminus\gamma_{t}$), as proven in [RS05]. The probability measure on these curves generated by the Brownian motion is called the $\mathrm{SLE}_{\kappa}$ measure in $\mathbb{H}$ from 00. Furthermore, $\gamma(t)\to\infty$ almost surely for every $\kappa>0$, which allows to define $\mathrm{SLE}_{\kappa}$ from $x$ to $y$ in $D$ a domain with $x\neq y\in\partial D$ as the image of $\mathrm{SLE}_{\kappa}$ in $\mathbb{H}$ (from 00 to $\infty$) by a conformal map $\varphi:\mathbb{H}\to D,\varphi(0)=x,\varphi(+\infty)=y$. We denote these measures $\nu_{D,\kappa}(z,w)$, probability measures for the moment. Because there remains freedom in precomposing $\varphi$ by a dilation, we will without further mention see $\nu_{D,\kappa}(z,w)$ as measures on curves up to (positive) time reparametrization.
The one-parameter family of measures $\left\{\nu_{D,\kappa}(z,w)\right\}_{(D,z,w)}$ in fact describes the only probability measures on curves from $z$ to $w$ in $D$ satisfying conformal invariance (clear from the construction) and a conformal Markov property, that states that the conditional law of $\gamma([t,+\infty))$ under $\nu_{D,\kappa}(z,w)$ given $\gamma_{t}$ is equal to $\nu_{D\setminus\gamma_{t},\kappa}(\gamma(t),w)$ (see [Sch00]). From now on we make the dependence on $\kappa$ implicit in the notation, and restrict to the range $\kappa\leq 4$ where the curves are simple.
In multiply connected domains, the two properties above are not sufficient to characterize $\mathrm{SLE}$ [^1] For instance, one could define a different SLE type of curve in a conformal annulus, that would behave like $\mathrm{SLE}_{2}$ when the conformal modulus is greater than one, like $\mathrm{SLE}_{4}$ otherwise. The conformal invariance requirement is too weak capture this change of behaviour of the curve because it cannot compare the law of the curve for different conformal classes of domains.. However, these curves satisfy an extra property, called the restriction property or boundary invariance rule [LSW03, Law09]: for $K\subset\mathbb{H}$ a compact hull such that $\mathrm{dist}(K,0)>0$ and $\mathbb{H}\setminus K$ is connected, it holds that $\nu_{\mathbb{H}\setminus K}(0,+\infty)\ll\nu_{\mathbb{H}}(0,+\infty)$, and that
(5) $\frac{\mathrm{d}\nu_{\mathbb{H}\setminus K}(0,+\infty)}{\mathrm{d}\nu_{\mathbb{H}}(0,+\infty)}(\eta)=\mathbf{1}_{\eta\subset\mathbb{H}\setminus K}\varphi^{\prime}(0)^{-h}\exp(\frac{\mathbf{c}}{2}m_{\mathbb{H}}(\eta,K)),$
where $\varphi:\mathbb{H}\setminus K\to\mathbb{H}$ is a conformal equivalence such that $\varphi(0)=0$ and $\varphi(z)=z+o(1)$ as $z\to\infty$. Notice that the term $\varphi^{\prime}(0)$ is positive and corresponds to the probability that a Brownian excursion from 00 to $\infty$ in $\mathbb{H}$ does not hit $K$ [LSW03]: similarly, for $D$ a simply connected (non-crossing) admissible domain and $D^{\prime}$ a test domain,
(6) $\frac{\mathrm{d}\nu_{D^{\prime}}(x,y)}{\mathrm{d}\nu_{D}(x,y)}(\eta)=\mathbf{1}_{\eta\subset D^{\prime}}\left(\frac{H_{\partial D^{\prime}}(x,y)}{H_{\partial D}(x,y)}\right)^{-h}\exp(\frac{\mathbf{c}}{2}m_{D}(\eta,D\setminus D^{\prime})),$
where $H_{\partial D^{\prime}}(x,y)/H_{\partial D}(x,y)$ is again the probability that an excursion from $x$ to $y$ in $D$ remains in $D^{\prime}$. It was suggested in [Law09] to use this extra property to characterize $\mathrm{SLE}$. Moreover, we now no longer see $\mathrm{SLE}$ as probability measures, but rather as measures with mass $Z_{D}(x,y)=Z_{D,\kappa}(x,y)$. Note that some regularity of $\partial D$ at $x$ and $y$ is required to define this mass or partition function: but this is not restrictive at the level of the corresponding probability measures, as any (finitely connected) domain is conformally equivalent to one with analytic boundary around the images of $x$ and $y$, so that the $\mathrm{SLE}$ probability measure can be defined there and brought back to the initial domain.
Definition 3.
$\mathrm{SLE}$ is the unique family of measures $\left\{\nu_{D}(x,y)\right\}_{(D,x,y)}$ on curves (modulo time-reparametrization) in $D$ a general admissible domain from $x$ to $y$ of total mass $Z_{D}(x,y)=\norm{\nu_{D}(x,y)}$ normalized so that $Z_{\mathbb{H}}(0,+\infty)=1$, satisfying the following:
  • Conformal covariance If $\varphi:D\to D^{\prime}$ is a conformal mapping, then the pushforward measure $\varphi_{\ast}\nu_{D}(x,y)$ satisfies
    (7) $\varphi_{\ast}\nu_{D}(x,y)=\absolutevalue{\varphi^{\prime}(x)}^{h}\absolutevalue{\varphi^{\prime}(y)}^{h}\nu_{D^{\prime}}(f(x),f(y)).$
    In particular, at the level of probability measures, denoting $\nu_{D}^{\#}(x,y):=\nu_{D}(x,y)/Z_{D}(x,y)$ if $Z_{D}(x,y)=\norm{\nu_{D}(x,y)}<+\infty$, conformal invariance holds:
    $\varphi_{\ast}\nu_{D}^{\#}(x,y)=\nu_{D^{\prime}}^{\#}(f(x),f(y)).$
  • Domain Markov property If $Z_{D}(x,y)<+\infty$, then the conditional probability measure of the remainder $\eta([t,+\infty))$ of a curve $\eta$ with respect to $\nu_{D}^{\#}(x,y)$ conditional on $\eta_{t}$ is given by $\nu_{D\setminus\eta_{t}}^{\#}(\eta(t),y)$.
  • Boundary perturbation Let $D^{\prime}\subset D$ be admissible domains agreeing in a neighborhood of $x$ and $y$. Then $\nu_{D^{\prime}}(x,y)\ll\nu_{D}(x,y)$, with Radon-Nikodym derivative given by
    (8) $\frac{\mathrm{d}\nu_{D^{\prime}}(x,y)}{\mathrm{d}\nu_{D}(x,y)}(\eta)=\mathbf{1}_{\eta\subset D^{\prime}}\exp(\frac{\mathbf{c}}{2}m_{D}(\eta,D\setminus D^{\prime})).$
Note that in the case of simply connected domains, it is consistent with Equations (5) and (6) to define $Z_{D}(x,y)=H_{D}(x,y)^{b}$ when $x,y\neq\infty$, and $Z_{D}(x,y)=\absolutevalue{\varphi^{\prime}(x)}^{b}$ when $y=\infty$, where $\varphi:D\to\mathbb{H}$ is a conformal equivalence with $\varphi(x)=0$ and $\varphi(z)=z+o(1)$ as $z\to\infty$.
In [Law11], these measures are constructed in any general admissible domain, from a Carathéodory argument: the restriction of $\nu_{D}(x,y)$ to curves contained in a general test domain $D^{\prime}$ is prescribed by the boundary perturbation rule, and the restriction to different test domains are compatible thanks to the additive structure of the loop measure. This justifies the denomination of ‘test domains’: to identify a measure on curves in a domain $D$ as the $\mathrm{SLE}$ measure it is sufficient to verify Equation (8) for $D^{\prime}\subset D$ simply connected. Note that uniqueness also follows from the construction.
We further stress that some conditions in the definition above are redundant: the domain Markov property for $\mathrm{SLE}$ in general domains is a consequence of the fact that this property holds on simply connected domains, and is inherited through the boundary perturbation rule (provided that the partition function is finite: see again [Law11] for all the details). Reversibility of the $\mathrm{SLE}$ is also inherited from the simply connected case, which was proven in [Zha08].
On the other hand, finiteness of the partition function is less clear. In [Law11], it is shown to be finite when $\kappa\leq 8/3$, or when $D$ is a conformal annulus. In this paper, we construct explicitly the $\mathrm{SLE}_{4}$ measure and the $\mathrm{SLE}_{\kappa}$ measure for $8/3<\kappa\leq 4$ in the non-crossing case, and provide an expression for their partition function that proves finiteness.
2. $\mathrm{SLE}_{4}$ as a level-line of Gaussian free fields
In this section we provide an explicit construction of the $\mathrm{SLE}$ measure in the specific case of $\kappa=4$, exploiting its relation with the Gaussian free field: this is possible both in the non-crossing and crossing case. First, we collect some classical facts about Gaussian free fields and $\mathrm{SLE}_{4}$.
2.1. Preliminaries
2.1.1. Gaussian free field
The (Dirichlet, with zero boundary conditions) Gaussian Free Field (GFF) in a domain $D$ is a Gaussian process $\Phi$ on $H_{0}^{1}(D)$, with mean zero and covariance function
$\mathbb{E}\left[\Phi_{f}\Phi_{g}\right]=\iint_{D\times D}f(x)G_{D}(x,y)g(y)\mathrm{d}x\mathrm{d}y<+\infty,$
where $G_{D}$ is the Green function, previously introduced. It is possible to define the GFF as a random element of $H^{0-}(D)$ tested against $H_{0}^{1}(D)$ (see for instance [WP21, Dub09]), so that instead of $\Phi_{f}$ we can write $(\Phi,f)$ to denote the distributional pairing. For our purposes it will suffice to see the GFF as an element of $H^{-1}(D)$. For $g$ a harmonic function in $D$, the GFF with boundary condition $g$ is understood as $\Phi^{g}:=\Phi^{0}+g$, where $\Phi^{0}$ is a zero-boundary GFF in $D$, and from now on we let $\Phi_{D}^{g}$ be our notation for a GFF in $D$ with mean $g$.
For $g\in H_{0}^{1}(D)$, $\Phi_{D}^{g}$ is also a zero-boundary GFF under the law that is also absolutely continuous to that of $\Phi_{D}^{0}$, with Radon-Nikodym derivative given by
$\exp((\Phi_{D}^{0},g)_{\nabla(D)}-(g,g)_{\nabla(D)}),$
where $(\cdot,\cdot)_{\nabla(D)}$ is the Dirichlet inner product and $(\Phi_{D}^{0},g)_{\nabla(D)}:=(\Phi_{D}^{0},-\Delta g)$.  [^2] Note that it is possible to naturally make sense of the GFF tested against functions in $H^{-1}(D)$, such as $-\Delta g\in H^{-1}(D)$: see [BP24]. This in fact only holds when $g\in H_{0}^{1}(D)$, identifying the latter as the Cameron-Martin space of the GFF.
The GFF also enjoys a natural domain Markov property. For $D$ a domain and $\delta$ a smooth crosscut splitting $D$ in $D_{1}$ and $D_{2}$, we can write
$\Phi_{D}^{0}=\Phi_{D_{1}}^{0}+u_{\Phi_{D}^{0}|_{\delta}}+\Phi_{D_{2}}^{0}$
where $u_{\Phi_{D}^{0}|_{\delta}}$ is a random harmonic function in $D\setminus\delta$, and all summands in the right-hand side are independent. In fact, one can define (see [Dub09]) a pathwise trace of $\Phi_{D}^{0}$ on $\delta$ as $\Phi_{D}^{0}|_{\delta}\in H^{-1}(\delta)$, whose law is a centered Gaussian process of covariance matrix $G_{D}|_{\delta\times\delta}$, so that $u_{\Phi_{D}^{0}|_{\delta}}$ is the harmonic extension of $\Phi_{D}^{0}|_{\delta}$ to $D\setminus\delta$ (which is indeed harmonic as the Poisson kernel is smoothing).
Note that the domain Markov property for the GFF holds in a much more general setting [WP21], although one may no longer interpret the random harmonic function $h$ as the extension of the values of the GFF on some compact set.
We are now interested in the study of random subsets of $D$, coupled with the GFF, for which a version of the Markov property holds.
Definition 4 (Local sets).
Given a triple $(\Phi=\Phi_{D},A,\Phi_{A})$ with $\Phi_{D}$ a GFF in $D$, $A\subset\overline{D}$ a random closed set and $\Phi_{A}$ a random distribution that can be seen as a harmonic function $h_{A}$ when restricted to $D\setminus A$, we say that $A$ is a local set for $\Phi$, or that $(\Phi,A,\Phi_{A})$ is a local set coupling, if conditionally on $(A,\Phi_{A})$, $\Phi-\Phi_{A}$ is a zero-boundary GFF in $D\setminus A$.
Changing the measure with a shift of the GFF given by some $g\in H_{0}^{1}(D)$ as in the Cameron-Martin theorem is also compatible with local sets, as shows the following lemma from [ALS22]:
Lemma 5 (Lemma 3.1 in [ALS22]).
Consider $(\Phi:=\Phi_{D}^{0},A,\Phi_{A})$ a local set coupling, $\tilde{\Phi}=\Phi-g$, and $g_{A}$ the orthogonal projection of $g$ in $H_{0}^{1}(D)$ on the subspace of functions that are harmonic in $D\setminus A$. Set $\tilde{\Phi}_{A}:=\Phi_{A}-g_{A}$. Then $(\tilde{\Phi},A,\tilde{\Phi}_{A})$ is a local set coupling and the law of $(A,\Phi_{A})$ under $\Phi$ is absolutely continuous with respect to the law of $(A,\tilde{\Phi}_{A}+g_{A})=(A,\Phi_{A})$ under $\tilde{\Phi}$, with Radon-Nikodym derivative given by
$\exp((\Phi_{A},g)_{\nabla}-\frac{1}{2}(g_{A},g_{A})_{\nabla}).$
Lastly, we will later need to compute the expectation of expressions similar to that appearing in the previous lemma, for $g$ a Gaussian random variable. The following lemma provides a formula that will be relevant in this computation.
Lemma 6 (Proposition 1.2.8 in [DPZ02]).
Let $H$ be a Hilbert space with $Q$ a positive, trace class operator, $M$ symmetric such that $Q^{1/2}MQ^{1/2}<I$, $m\in H$. Denote by $N_{Q}$ the centered Gaussian measure on $H$ with covariance $Q$. Then
$\mathbb{E}^{N_{Q}}\left[\exp(\frac{1}{2}\left\langle Mh,h\right\rangle_{H}+\left\langle m,h\right\rangle_{H})\right]\quad$
$=\left(\mathrm{det}_{F}(1-Q^{1/2}MQ^{1/2})\right)^{-1/2}$ $\exp(\frac{1}{2}\norm{(1-Q^{1/2}MQ^{1/2})^{-1/2}Q^{1/2}m}_{H}^{2}).$
2.1.2. Coupling between $\mathrm{SLE}_{4}$ and the GFF
In simply connected domains, there is already a deep connection between the GFF and $\mathrm{SLE}_{4}$ [SS13, Dub09, WP21].
Theorem 7 ($\mathrm{SLE}_{4}$ is a local set for the GFF).
Let $D$ be a simply connected domain with $x\neq y\in\partial D$. There exists a coupling $(\Phi_{D}^{g_{D}},\eta,\Phi_{\eta})$ with $\eta\sim\nu_{D,4}^{\#}(x,y)$ an $\mathrm{SLE}_{4}$ curve from $x$ to $y$ in $D$ that is a local set coupling, with $\Phi_{\eta}$ given by $h_{\eta}$ on $D\setminus\eta$, and $h_{\eta}(z):=-\lambda+2\lambda\mathbf{1}(z\in(D\setminus\eta)_{\text{right}})$.
In other words, there is a local set coupling $(\Phi_{D}^{g_{D}},\eta,h_{\eta})$ in which the marginal law of $\eta$ is that of an $\mathrm{SLE}_{4}$ from $x$ to $y$ and such that conditionally on $\eta$, $\Phi_{D}^{g_{D}}=\Phi_{D\setminus\eta,\text{left}}^{-\lambda}+\Phi_{D\setminus\eta,\text{right}}^{+\lambda}$, a sum of two independent GFFs on each side of $D\setminus\eta$. When the domain is no longer simply connected, or the harmonic function $g_{D}$ is replaced with a bounded harmonic function with different boundary conditions away from the target points $x$ and $y$, it is still possible to make sense of a level-line of the GFF.
Proposition 8 (General level-lines, [ASW19, Lemma 15]).
Let $D$ be a non-crossing admissible domain and $g$ a bounded harmonic function in $D$ with Dirichlet boundary conditions $-\lambda$ (resp. $+\lambda$) on $(yx)$ (resp. $(xy)$), arbitrary on other parts of $\partial D$. Then there exists a unique (up to time-reparametrization) law on random simple curves $(\eta(t),0\leq t\leq\tau_{D})$ in $D$ with $\eta(0)=x$ that can be coupled with a GFF $\Phi_{D}^{g}$ such that $(\Phi_{D}^{g},\eta,\Phi_{\eta})$ is a local set coupling, and $\Phi_{\eta}$ corresponds to the unique harmonic function in $D\setminus\eta$ with boundary values given by $g$ on $\partial D$ and $-\lambda$ (resp. $+\lambda$) on the left side of $\eta_{\tau_{D}}$ (resp. on the right side of $\eta_{\tau_{D}}$). We call $\eta$ the generalized level-line of $\Phi_{D}^{g}$.
It is also worth mentioning that the probability that $\tau_{D}<+\infty$ and that $\eta(t)$ converges to a point in $\partial D\setminus\left\{y\right\}$ as $t\to\tau_{D}^{-}$ is equal to 00 if and only if |$\absolutevalue{g}\geq\lambda$ on $\partial D$ [PW17, ASW19]: in this case, the level-line can be parametrized for instance by its capacity in $\mathbb{D}$ seen from $y$, so that $\eta(t)\xrightarrow[t\to\infty]{}y$ a.s. Furthermore, in this coupling $\eta$ is measurable with respect to $\Phi_{D}^{g}$.
In the case of $D$ a crossing domain, generalized level-lines can also be defined for GFFs $\Phi_{D}^{g}$, where this time $g$ is multivalued with boundary values locally constant with a change of ±$\pm 2\lambda$ around $x$ and $y$ on $\partial D$ (see Section 2.3). This was proven for instance in [IK13] for the case of conformal annuli, and extends to the general case by absolute continuity. For the sake of completeness, we include a more thorough explanation in Appendix A.
2.2. Constructing the non-crossing $\mathrm{SLE}_{4}$ measure
In this section, we construct a measure on curves joining two points in the same boundary component of an $n$-connected domain $D$ (non-crossing case). We assume that no boundary component is reduced to a point. Such domains are conformally equivalent to a non-crossing admissible domain as defined in Section 1.1: we can actually assume that the boundary components different to $\partial\mathbb{D}$ are circles, in particular they are smooth. Because the measure we construct is conformally invariant, we restrict the study to this kind of domain.
For $\mathbf{b}=(b_{1},\dots,b_{n})\in\left\{\pm\lambda\right\}^{n}$, we define the event:
$\mathcal{A}_{\mathbf{b}}:=\left\{\eta\text{ curve from }x\text{ to }y\text{ in}D\,|\,\forall i:b_{i}=-\lambda,\mathcal{B}_{i}\subset(D\setminus\eta)_{\text{left}}\text{ and }\forall i:b_{i}=+\lambda,\mathcal{B}_{i}\subset(D\setminus\eta)_{\text{right}}\right\},$
i.e. the set of curves from $x$ to $y$ in $D$ leaving the components of $\mathcal{B}$ corresponding to the positive components of $\mathbf{b}$ to the left, the others to the right. Note that these events partition $\left\{\eta\subset D\right\}$, since $\mathcal{A}_{\mathbf{b}}\cap\mathcal{A}_{\mathbf{b}^{\prime}}=\emptyset$ for $\mathbf{b}^{\prime}\neq\mathbf{b}$, and that any simple curve from $x$ to $y$ in $D$ must belong to $\mathcal{A}_{\mathbf{b}^{\prime}}$ for some $\mathbf{b}^{\prime}\in\left\{\pm\lambda\right\}^{n}$.
Let also $\mathcal{B}_{\mathbf{b}}^{-}$ (resp. $\mathcal{B}_{\mathbf{b}}^{+}$) correspond to the union of $\mathcal{B}_{l}$ (resp. $\mathcal{B}_{r}$) and of the boundary components $\mathcal{B}_{i}$ such that $b_{i}=-\lambda$ (resp. $b_{i}=+\lambda$).
Figure 1. An example of a domain with the definitions above, together with a test domain $D^{\prime}$ shaded in light blue. Here the curve $\eta$ depicted in the middle is in $\mathcal{A}_{\mathbf{b}}$ for $\mathbf{b}=\left\{-\lambda,+\lambda,+\lambda\right\}$. In particular, the boundary values of $g_{\mathbf{b}}$ are $-\lambda$ and $+\lambda$ exactly on the parts of $\partial D$ that lie the left and on the right of $\eta$, respectively.
Figure 2. An example of two non path-homotopic curves in the event $\mathcal{A}_{\left\{1\right\}}$. In general, there are infinitely (countably) many homotopy classes of distinct curves corresponding to every event $\mathcal{A}_{\mathbf{b}}$.
Note that the events $\mathcal{A}_{\mathbf{b}}$ do not quite fix the homotopy class of the curve, see Figure 2. However, for $D^{\prime}$ a test domain relative to $D$, there is a unique $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$ such that $\left\{\eta\in D^{\prime}\right\}\subset\left\{\eta\in\mathcal{A}_{\mathbf{b}}\right\}$.
We define $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ as the law of the generalized level line of a GFF $\Phi_{D}$ in $D$ with boundary conditions given by $g_{\mathbf{b}}$, the harmonic function with boundary values $-\lambda$ on $\mathcal{B}_{\mathbf{b}}^{-}$, $+\lambda$ on $\mathcal{B}_{\mathbf{b}}^{+}$. The important picture to keep in mind is that on the event $\mathcal{A}_{\mathbf{b}}$ and conditionally on $\eta$ the level-line of the GFF, the latter is equal in law to a GFF in $D\setminus\eta$ with piecewise constant mean: $-\lambda$ on the left of $\eta$, $+\lambda$ on its right. That would not be true if $\eta\notin\mathcal{A}_{\mathbf{b}}$.
Theorem 9.
Let $D\subset\mathbb{D}$ be admissible. The measure $\nu_{D}$ on simple curves from $x$ to $y$ in $D$ defined as
(9) $\nu_{D}(\eta):=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\mathbb{P}_{D}^{g_{\mathbf{b}}}(\eta)\cdot Z_{D,\mathbf{b}},$
where
(10) $Z_{D,\mathbf{b}}=\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2}),$
is the measure of $\mathrm{SLE}_{4}$ in $(D,x,y)$.
Remark 2.1.
There are several equivalent ways of expressing the mass term $Z_{D,\mathbf{b}}$ (see also Section 2.2). It will appear in the proof that for $D^{\prime\prime}\subset D$ a (non-crossing) test domain, one has
(11) $\frac{Z_{D,\mathbf{b}}}{Z_{D^{\prime\prime}}}$ $=\exp(\frac{1}{2}(\norm{g_{D^{\prime\prime}}}_{\nabla(D^{\prime\prime})}^{2}-\norm{g_{\mathbf{b}}}_{\nabla(D)}^{2}))$
well-defined without the need for regularization. Here again this might be rewritten
$\frac{Z_{D,\mathbf{b}}}{Z_{D^{\prime\prime}}}=\exp(2\lambda^{2}(M(\mathcal{B}_{l}^{\prime\prime}\cup\mathcal{B}_{l},\mathcal{B}_{r}^{\prime\prime}\cup\mathcal{B}_{r})-M(\mathcal{B}_{\mathbf{b}}^{-},\mathcal{B}_{\mathbf{b}}^{+}))),$
in terms of inverse extremal distance, see Section 1.3.
This is consistent with Equation (10), given that $Z_{D^{\prime\prime}}=H_{D^{\prime\prime}}(x,y)^{h}$ and that in [Dub09] it is argued (proof of Proposition 5.2, page 29) that
$\exp(-\frac{1}{2}\norm{g_{D^{\prime\prime}}}_{\nabla(D^{\prime\prime}),\text{reg}}^{2})=\mathrm{const}\cdot H_{D^{\prime\prime}}(x,y)^{\nicefrac{{1}}{{4}}},$
where $\nicefrac{{1}}{{4}}$ is the boundary scaling exponent $h_{\kappa}$ for $\kappa=4$.
Before going to the proof of the Theorem, we make some simple observations about a GFF in $D$ and its coupling with a level-line. We fix $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$. Let $(\Phi_{D},\eta)$ be a local set with $\Phi_{D}:=\Phi_{D}^{g_{\mathbf{b}}}$ a GFF in $D$ with boundary values given by $g_{\mathbf{b}}$. On the event $\mathcal{A}_{\mathbf{b}}$, the harmonic part $h_{\eta}$ is two-valued and reads
(12) $h_{\eta}(z)=(-\lambda+2\lambda\mathbf{1}_{\text{ right of }\eta}(z)).$
In other words, conditionally on $\eta$, $\Phi_{D}$ is equal in law to the sum of two independent GFFs $\Phi_{(D\setminus\eta)_{\text{left}}}^{0}-\lambda,\Phi_{(D\setminus\eta)_{\text{right}}}^{0}+\lambda$ on the left (resp. right) side of $\eta$ in $D$. We write, conditionally on $\eta$ and on the event $\mathcal{A}_{\mathbf{b}}$,
$\Phi_{D}=\Phi_{D\setminus\eta}^{0}+h_{\eta},$
where $\Phi_{D\setminus\eta}^{0}$ should be understood as $\Phi_{(D\setminus\eta)_{\text{left}}}^{0}+\Phi_{(D\setminus\eta)_{\text{right}}}^{0}$. Let us now consider $D^{\prime}\subset D$ a test domain with $\left\{\eta\subset D^{\prime}\right\}\subset\mathcal{A}_{\mathbf{b}}$. Using the standard Markov property on $\Phi_{D\setminus\eta}^{0}$ on the event that $\eta\subset D^{\prime}$, we can further expand:
$\Phi_{D}|_{D^{\prime}}=\Phi_{D^{\prime}\setminus\eta}^{0}+u_{\mathcal{B}_{l}^{\prime}\cup\mathcal{B}_{r}^{\prime}}^{D^{\prime}\setminus\eta}+h_{\eta}|_{D^{\prime}},$
where $u_{\mathcal{B}^{\prime}}^{D^{\prime}\setminus\eta}$ is a harmonic function in $D^{\prime}\setminus\eta$ corresponding to the harmonic extension of the restriction $\xi_{l}$ (resp. $\xi_{r}$) of $\Phi_{D\setminus\eta}^{0}$ on $\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}^{\prime}$), and of zero on the other parts of the boundary (including $\eta$). In other words, since $\xi=(\xi_{l},\xi_{r})$ is distributed like a centered Gaussian process on $\mathcal{B}_{l}^{\prime}\cup\mathcal{B}_{r}^{\prime}$ whose covariance matrix is diagonal with elements $(G_{D\setminus\eta})|_{\mathcal{B}_{l}^{\prime}}$ and $(G_{D\setminus\eta})|_{\mathcal{B}_{r}^{\prime}}$, conditionally on $\eta$ still and on the event $\left\{\eta\subset D^{\prime}\right\}\subset\mathcal{A}_{\mathbf{b}}$, we can write $\Phi_{D}|_{D^{\prime}}$ as $\Phi_{D^{\prime}\setminus\eta}^{0}+\bar{h}_{\eta}$, with $\bar{h}_{\eta}$ is a random harmonic function in $D^{\prime}\setminus\eta$ with boundary values
$\bar{h}_{\eta}=\begin{dcases}-\lambda&:\text{ on the left-side of }\eta\\ +\lambda&:\text{ on the right-side of }\eta\\ \xi_{l}\sim\mathcal{N}(-\lambda,G_{D\setminus\eta}|_{\mathcal{B}_{l}^{\prime}})&:\text{ on }\mathcal{B}_{l}^{\prime}\\ \xi_{r}\sim\mathcal{N}(+\lambda,G_{D\setminus\eta}|_{\mathcal{B}_{r}^{\prime}})&:\text{ on }\mathcal{B}_{r}^{\prime}\end{dcases}.$
We consider, similarly, another local set $(\tilde{\Phi}_{D^{\prime}},\tilde{\eta})$ in $D^{\prime}$ with $\tilde{\Phi}_{D^{\prime}}=\tilde{\Phi}_{D^{\prime}}^{g_{D^{\prime}}}$, so that conditionally on $\tilde{\eta}$ we can write $\tilde{\Phi}_{D^{\prime}}=\tilde{\Phi}_{D^{\prime}\setminus\tilde{\eta}}^{0}+h_{\tilde{\eta}}$, with
$h_{\tilde{\eta}}(z)=(-\lambda+2\lambda\mathbf{1}_{\text{ right of }\tilde{\eta}}(z)).$
For computations, it will sometimes be easier to work with a GFF with zero-boundary conditions, coupled with a level-line such that the resulting harmonic part $h_{\tilde{\eta}}$ is no longer only a two-valued function, but is subtracted $g_{D^{\prime}}$. The marginal law of $\tilde{\eta}$ in this coupling will be denoted $\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}$ (the dependency in $x$,$y$ is implicit), and in this case is equal to the $\mathrm{SLE}_{4}$ measure $\nu_{D}^{\#}(z,w)$ by Theorem 7 (from now on we will only be concerned with $\mathrm{SLE}_{4}$ and therefore let $\nu_{D}$ implicitly refer to this value of $\kappa$). On the other hand, the marginal law of $\eta$ in the coupling $(\eta,\Phi_{D})$ as presented above on the event $\left\{\eta\subset D^{\prime}\right\}$ and given the values of $\xi$ is denoted $\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}$. It therefore corresponds to the law of the level-line of a GFF in $D^{\prime}$ whose mean is the harmonic extension of $\xi$ on $\mathcal{B}_{l}^{\prime}\cup\mathcal{B}_{r}^{\prime}$ and $-\lambda$ (resp. $+\lambda$) on $\mathcal{B}_{l}\setminus\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}\setminus\mathcal{B}_{r}^{\prime}$).
We now turn to the proof of the Theorem, that goes in three steps: we first establish a purely measure-theoretic Lemma, which then allows us to compute the Radon-Nikodym derivative by fixing $\xi$, and then averaging. The two latter steps are done in a Proposition and a Lemma.
Lemma 10.
Let $X,Y_{1},Y_{2}$ be random variables on the same probability space, taking values in a Polish space $E$, with associated distributions $\mathbb{P}_{X},\mathbb{P}_{Y_{1}},\mathbb{P}_{Y_{2}}$ respectively. Denote also by $\mathbb{P}_{Y_{1}|Y_{2}=y_{2}}$ the conditional law of $Y_{1}$ on ${Y_{2}=y_{2}}$ (similarly for $\mathbb{P}_{Y_{2}|Y_{1}=x}$) and assume that $\mathbb{P}_{X}\ll\mathbb{P}_{Y_{1}|Y_{2}=y_{2}}$, for $\mathbb{P}_{Y_{2}}$-almost every $y_{2}\in E$. The following equality between Radon-Nikodym derivatives holds $\mathbb{P}_{X}$-a.s.
$\frac{\mathrm{d}\mathbb{P}_{X}}{\mathrm{d}\mathbb{P}_{Y_{1}}}(x)=\mathbb{E}_{Y_{2}|Y_{1}=x}\left[\frac{\mathrm{d}\mathbb{P}_{X}}{\mathrm{d}\mathbb{P}_{Y_{1}|Y_{2}=y}}(x)\bigg{|}_{y=Y_{2}}\right].$
Proof.
See Appendix B. ∎
Proof of Theorem 9.
We proceed by induction on $n$, the number of connected components of $\mathcal{B}$ that do not contain $x$ and $y$. For clarity we temporarily denote the measures in the statement of the theorem as $\tilde{\nu}_{D}$, as we are yet to prove that they correspond to the $\mathrm{SLE}_{4}$ measures.
When $n=0$, Equation (9) defines $\tilde{\nu}_{D}$ as the law of the level line of a GFF in $D$ with boundary conditions $-\lambda$ on the counterclockwise arc $y\to x$, $+\lambda$ on the other, which corresponds to $\mathrm{SLE}_{4}$ in $D$ upon setting $Z_{D}=H_{D}(x,y)^{\nicefrac{{1}}{{4}}}$.
Let now $n\geq 1$ and $D^{\prime\prime}\subset D$ be a test domain: because $D^{\prime\prime}$ is simply connected by assumption, there exists $\mathbf{b}^{\prime}=(b_{1}^{\prime},\dots,b_{n}^{\prime})\in\left\{\pm\lambda\right\}^{n}$ such that $\left\{\eta\subset D^{\prime\prime}\right\}\subset\mathcal{A}_{\mathbf{b}^{\prime}}$. Up to relabeling the components of $\mathcal{B}$ we may assume that $\mathbf{b}^{\prime}=\left\{-\lambda,\dots,-\lambda,+\lambda,\dots,+\lambda\right\}$ with $1\leq k\leq n$ such that $b_{k}^{\prime}=-\lambda,b_{k+1}^{\prime}=+\lambda$, so that if $D^{\prime}$ is the union of $D^{\prime\prime}$ and the component of $D\setminus D^{\prime\prime}$ on the left of $\eta$, we have by induction that
$\frac{\mathrm{d}\nu_{D^{\prime\prime}}}{\mathrm{d}\tilde{\nu}_{D^{\prime}}}(\eta)=\mathbf{1}_{\eta\subset D^{\prime\prime}}\exp(\frac{1}{2}m_{D^{\prime}}(\eta,D^{\prime}\setminus D^{\prime\prime})).$
Figure 3.
We have to show that $\tilde{\nu}_{D^{\prime}}$ is absolutely continuous w.r.t. $\tilde{\nu}_{D}$, with Radon-Nikodym derivative given by
(13) $\frac{\mathrm{d}\tilde{\nu}_{D^{\prime}}}{\mathrm{d}\tilde{\nu}_{D}}(\eta)=\sum_{\begin{subarray}{c}\mathbf{b}=(\mathbf{b}_{k},+\lambda,\dots,+\lambda)\\ \mathbf{b}_{k}\in\left\{\pm\lambda\right\}^{k}\end{subarray}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}_{k}}}}{\mathrm{d}\mathbb{P}_{D}^{g_{\mathbf{b}}}}(\eta)\frac{Z_{D^{\prime},\mathbf{b}_{k}}}{Z_{D,\mathbf{b}}}=\sum_{\begin{subarray}{c}\mathbf{b}=(\mathbf{b}_{k},+\lambda,\dots,+\lambda)\\ \mathbf{b}_{k}\in\left\{\pm\lambda\right\}^{k}\end{subarray}}\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}\exp(\frac{1}{2}m_{D}(\eta,D\setminus D^{\prime})).$
Indeed, by the additive property of the loop measure we would then obtain
$\frac{\mathrm{d}\nu_{D^{\prime\prime}}}{\mathrm{d}\tilde{\nu}_{D}}(\eta)=\frac{\mathrm{d}\nu_{D^{\prime\prime}}}{\mathrm{d}\tilde{\nu}_{D^{\prime}}}(\eta)\frac{\mathrm{d}\tilde{\nu}_{D^{\prime}}}{\mathrm{d}\tilde{\nu}_{D}}(\eta)$ $=\mathbf{1}_{\eta\subset D^{\prime}}\exp(\frac{1}{2}m_{D^{\prime}}(\eta,D^{\prime}\setminus D^{\prime\prime})+\frac{1}{2}m_{D}(\eta,D\setminus D^{\prime}))$
$=\mathbf{1}_{\eta\subset D^{\prime}}\exp(\frac{1}{2}m_{D}(\eta,D\setminus D^{\prime\prime})).$
To carry out this computation, we use Lemma 10 with $\mathbb{P}_{X}=\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}$ (now abusing notation to drop the superscript $\mathbf{b}_{k}$), $\mathbb{P}_{Y_{1}}=\mathbb{P}_{D}^{g_{\mathbf{b}}}$, and $\mathbb{P}_{Y_{2}}$ corresponding to the law of $\xi:=T_{\mathcal{B}_{r}^{\prime}}\Phi_{D}^{0}$, the restriction of the GFF to $\mathcal{B}_{r}^{\prime}$ (recall that $\mathcal{B}_{r}^{\prime}=(\mathcal{B}^{\prime}\setminus\mathcal{B})\cap D$). As noted earlier, conditionally on $\eta\sim\mathbb{P}_{D}^{g_{\mathbf{b}}}$ and $\left\{\eta\subset D^{\prime}\right\}$, we have that $\xi\sim\mathcal{N}(0,G_{D\setminus\eta}|_{\mathcal{B}_{r}^{\prime}})$. We can then write
$\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}}{\mathrm{d}\mathbb{P}_{D}^{g_{\mathbf{b}}}}=\mathbb{E}^{\mathcal{N}(0,G_{D\setminus\eta}|_{\mathcal{B}_{r}^{\prime}})}\left[\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}}{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}}(\eta)\right].$
The fact that the term in the expectation is well-defined and its precise value are provided by the following proposition:
Proposition 11.
For all $\xi\in H^{-1}(\mathcal{B}_{r}^{\prime})$, it holds that $\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}\ll\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}$, and we have explicitly that
(14) $\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}}{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}}(\eta)=\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\in D^{\prime}}\exp(\frac{1}{2}(\xi,(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})\xi)_{\mathcal{B}_{r}^{\prime}}-2\lambda((H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}},\xi)_{\mathcal{B}_{r}^{\prime}}).$
We first pursue the proof of the theorem with this proposition in hand, and will turn to its proof just after. The computation of the expectation now relies on applying Lemma 6 to Equation (14).
To simplify the corresponding expression, first notice that, thanks to Lemma 1,
$I-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})G_{D\setminus\eta}=I-H_{D^{\prime}}H_{D\setminus\eta}.$
We dropped the subscripts and superscripts indicating that the operators are seen on $H^{-1}(\mathcal{B}_{r}^{\prime})$, for more readability. Consequently, the determinant term in the expectation of Equation (14) simplifies to
$(\mathrm{det}_{F}(I-(G_{D\setminus\eta})^{1/2}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})(G_{D\setminus\eta})^{1/2}))^{-1/2}$ $=(\mathrm{det}_{F}(I-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})(G_{D\setminus\eta})))^{-1/2}$
$=\exp(\frac{1}{2}m_{D}(\eta,\mathcal{B}_{r}^{\prime}))$
by Proposition 2 (note that this Proposition implicitly contains the statement that the determinant is well-defined, and hence that the use of Lemma 6 is legitimate).
On the other hand, the term corresponding to the exponential in Lemma 6 can be rewritten as
$\exp(\frac{1}{2}\norm{2\lambda(I-G_{D\setminus\eta}^{1/2}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})G_{D^{\prime}\setminus\eta}^{1/2})^{-1/2}G_{D\setminus\eta}^{1/2}(H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}}}^{2})$
$=$ $\exp(2\lambda^{2}((H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}},G_{D\setminus\eta}(I-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})G_{D\setminus\eta})^{-1}(H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}})_{\mathcal{B}_{r}^{\prime}})$
$=$ $\exp(2\lambda^{2}((H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}},G_{D}(H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}})_{\mathcal{B}_{r}^{\prime}})$
(15) $=$ $\exp(2\lambda^{2}(1_{\mathcal{B}_{\mathbf{b}}^{-}},(H_{\partial D}-H_{\partial D^{\prime}})1_{\mathcal{B}_{\mathbf{b}}^{-}})_{\mathcal{B}_{\mathbf{b}}^{-}}).$
This can be re-written as a difference of Dirichlet energies, as in Section 1.3, so that we finally obtain
$\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}}{\mathrm{d}\mathbb{P}_{D}^{g_{\mathbf{b}}}}=\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}\exp(\frac{1}{2}m_{D}(\eta,D\setminus D^{\prime}))\exp(\frac{1}{2}(\norm{g_{D,\mathbf{b}}}_{\nabla(D)}^{2}-\norm{g_{D^{\prime},\mathbf{b}}}_{\nabla(D^{\prime})}^{2})).$
Replacing this in Equation (13) with the right expression of $Z_{D,\mathbf{b}},Z_{D^{\prime},\mathbf{b}}$ finishes the proof. ∎
Let us mention as a remark that the reason we rely on induction is to avoid having to deal with both boundaries at once. It would be possible to consider $D^{\prime}$ a test domain, $\xi=(\xi_{l},\xi_{r})$ a Gaussian process on $\mathcal{B}_{l}^{\prime}\times\mathcal{B}_{r}^{\prime}$ with covariance matrix having diagonal values $(G_{D})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}},(G_{D})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}$, and follow the same approach as above: however the expression (14) and the computation of the Gaussian integral that results become more involved.
We now turn to the proof of Proposition 11.
Proof of Proposition 11.
Consider $D^{\prime}_{\varepsilon}:=\left\{x\in D^{\prime}:\mathrm{dist}(x,\mathcal{B}_{r}^{\prime})>\varepsilon\right\}$, $\mathcal{B}^{\prime}_{\varepsilon}:=\partial D^{\prime}_{\varepsilon}\setminus\partial D^{\prime}$ and define $u_{\xi}^{\varepsilon}$ as the unique harmonic function continuous on $\overline{D^{\prime}}$ and harmonic on $D^{\prime}\setminus\mathcal{B}^{\prime}_{\varepsilon}$ with boundary values
$u_{\xi}^{\varepsilon}=\begin{cases}u_{\xi}&\text{on }\mathcal{B}^{\prime}_{\varepsilon}\\ 0&\text{on }\partial D^{\prime},\end{cases}$
where $u_{\xi}$ is the harmonic extension of $\xi$ to $D^{\prime}$. Note that $u_{\xi}^{\varepsilon}$ and $u_{\xi}$ coincide in $D^{\prime}_{\varepsilon}$. Let $\mathbb{Q}$ be the law of $\Phi:=\Phi_{D^{\prime}}^{0}+g_{\mathbf{b}}$ and ~$\tilde{\mathbb{Q}}_{\varepsilon}$ defined as
(16) $\frac{\mathrm{d}\tilde{\mathbb{Q}}_{\varepsilon}}{\mathrm{d}\mathbb{Q}}=\exp((\Phi,u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}-\frac{1}{2}(u_{\xi}^{\varepsilon},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})})$
so that $\tilde{\Phi}:=\Phi-u_{\xi}^{\varepsilon}$ is a zero-boundary GFF in $D^{\prime}$ under ~$\tilde{\mathbb{Q}}_{\varepsilon}$, by Girsanov. Let $\eta,\tilde{\eta}$ be the level-lines for $\Phi,\tilde{\Phi}+u_{\xi}$ respectively.
Like in [ALS22, Proposition 3.3], we argue that the events $\left\{\mathrm{dist}(\eta,\mathcal{B}_{r}^{\prime})>\varepsilon\right\},\left\{\mathrm{dist}(\tilde{\eta}_{\varepsilon},\mathcal{B}_{r}^{\prime})>\varepsilon\right\}$ are (~$\tilde{\mathbb{Q}}_{\varepsilon}$-)a.s. equal, and that on these events $\eta$ and $\tilde{\eta}_{\varepsilon}$ coincide. Whereas $(\eta,\Phi_{\eta})$ is a local set under $\mathbb{Q}$ by definition, $(\eta_{\tau},\Phi_{\eta_{\tau}})$ is also a local set [SS09, Lemma 3.12], where $\tau$ is the exit time of $\eta$ of $D_{\varepsilon}^{\prime}$, and $\Phi_{\eta_{\tau}}$ is the harmonic function in $D^{\prime}\setminus\eta_{\tau}$ with boundary values given by $-\lambda$ (resp. $+\lambda$) on the left (resp. right) boundary of $\eta_{\tau}$, and coinciding with $g_{\mathbf{b}}$ on the other parts of $\partial D^{\prime}$. Now by Lemma 5, $(\eta_{\tau},\tilde{\Phi}_{\eta_{\tau}})$ is also a local set under ~$\tilde{\mathbb{Q}}_{\varepsilon}$, with $\tilde{\Phi}_{\eta_{\tau}}=\Phi_{\eta_{\tau}}-(u_{\xi}^{\varepsilon})_{\eta_{\tau}}$. On the other hand, $(\tilde{\eta}_{\tau},\tilde{\Phi}_{\eta_{\tau}})$, the level-line for $\tilde{\Phi}+u_{\xi}$, is also a local set under ~$\tilde{\mathbb{Q}}_{\varepsilon}$. Furthermore, the fields $\Phi$ and $\tilde{\Phi}+u_{\xi}$ coincide on $D_{\varepsilon}^{\prime}$. It follows from local uniqueness of level-lines ([ASW19, Lemma 15]) that $\eta_{\tau}=\tilde{\eta}_{\tau}$$\tilde{\mathbb{Q}}_{\varepsilon}$-almost surely. If $\tau=+\infty$, i.e. $\eta\subset D_{\varepsilon}^{\prime}$, then it is clear that $\tilde{\eta}\subset D_{\varepsilon}^{\prime}$ and $\eta=\tilde{\eta}$.
This implies that the law $\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}$ of the level line of $\Phi+u_{\xi}$ under ~$\tilde{\mathbb{Q}}_{\varepsilon}$, on the event that it stays at a distance $>\varepsilon$ from $\mathcal{B}_{r}^{\prime}$, is absolutely continuous w.r.t. the law $\mathbb{P}_{D^{\prime}}$ of the level line of $\Phi$ under $\mathbb{Q}$ on the event that it too stays at a distance $>\varepsilon$ from $\mathcal{B}_{r}^{\prime}$, and the Radon-Nikodym derivative is given the conditional expectation of the RHS in Equation (16) given $\eta$, which by Lemma 5 is given by:
(17) $\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}}}}{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{\mathbf{b}},\xi}}(\eta)=\exp(-(\Phi_{\eta},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}+\frac{1}{2}((u_{\xi}^{\varepsilon})_{\eta},(u_{\xi}^{\varepsilon})_{\eta})_{\nabla(D^{\prime})}).$
We stress here that the Lemma above applies to a zero-boundary GFF, and a local set with $\Phi_{\eta}=h_{\eta}=0$ on the boundary. In our framework, that amounts to considering the curve $\eta$ coupled with $\Phi-g_{\mathbf{b}}=\Phi_{D^{\prime}}^{0}$, and where the harmonic part $h_{\eta}$ is the difference of the two-valued function that is equal to $-\lambda$ on the left of $\eta$ and $+\lambda$ on the right, and of the harmonic function $g_{\mathbf{b}}$. We compute this expression on the event that $\mathrm{dist}(\eta,\mathcal{B}_{r}^{\prime})>\varepsilon$, and take $\varepsilon$ to 00. Since $u_{\xi}$ is harmonic in $D^{\prime}$, we can write
$(\Phi_{\eta},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}$ $=(\Phi_{\eta},-\Delta u_{\xi}^{\varepsilon})_{D^{\prime}}+\int_{\partial D^{\prime}}\Phi_{\eta}\cdot\partial_{\nu}u_{\xi}^{\varepsilon}$
$=(\Phi_{\eta},-\Delta(u_{\xi}^{\varepsilon}-u_{\xi}))_{D^{\prime}}+\int_{\partial D^{\prime}}h_{\eta}\cdot\partial_{\nu}u_{\xi}^{\varepsilon}$
$=-\int_{\partial D^{\prime}}\partial_{\nu}(u_{\xi}^{\varepsilon}-u_{\xi})\cdot h_{\eta}+\int_{\partial D^{\prime}}(u_{\xi}^{\varepsilon}-u_{\xi})\cdot\partial_{\nu}h_{\eta}+\int_{\partial D^{\prime}}h_{\eta}\cdot\partial_{\nu}u_{\xi}^{\varepsilon}$
(18) $=-(\xi,\partial_{\nu}h_{\eta})_{\mathcal{B}_{r}^{\prime}}$
using that $h_{\eta}$ is harmonic and zero on $\partial D^{\prime}$, and $u_{\xi}^{\varepsilon}-u_{\xi}$ vanishes on $\partial(D^{\prime}\setminus D^{\prime}_{\varepsilon})$. Note again the abuse of notation: the above denotes the distributional pairing between $\xi\in H^{-1}(\mathcal{B}_{r}^{\prime})$ and $\partial_{\nu}h_{\eta}\in H^{1}(\mathcal{B}_{r}^{\prime})$. The explicit expression for $h_{\eta}$ (or for $g_{\mathbf{b}}$) in terms of the Poisson kernel allows to further write
$(\Phi_{\eta},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}=\left(\xi,2\lambda\int_{\mathcal{B}_{\mathbf{b}}^{-}}H_{\partial D^{\prime}}(\mathrm{d}y,\cdot)\right)_{\mathcal{B}_{r}^{\prime}}=2\lambda(\xi,(H_{\partial D^{\prime}})_{\mathcal{B}_{\mathbf{b}}^{-}}^{\mathcal{B}_{r}^{\prime}}1_{\mathcal{B}_{\mathbf{b}}^{-}})_{\mathcal{B}_{r}^{\prime}}.$
On the other hand, we claim that
(19) $((u_{\xi}^{\varepsilon})_{\eta},(u_{\xi}^{\varepsilon})_{\eta})_{\nabla(D^{\prime})}=(\xi,(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})\xi)_{\mathcal{B}_{r}^{\prime}}$
Notice that this expression makes sense as $H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)}$ is a smoothing operator, and that it suffices to prove it for $\xi$ smooth as for a general $\xi\in H^{-1}(\mathcal{B}_{r}^{\prime})$, taking $(\xi_{n})_{n\geq 0}\subset C_{0}^{\infty}(\mathcal{B}_{r}^{\prime})$ converging to $\xi$ in $H^{-1}(\mathcal{B}_{r}^{\prime})$, we can write Equation (19) for $\xi_{n}$ and take the limit $n\to\infty$. This follows from $H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)}$ being continuous on $H^{-1}(D^{\prime})$ and by $(u_{\xi_{n}}^{\varepsilon})_{\eta}$ converging to $(u_{\xi}^{\varepsilon})_{\eta}$ in $H_{0}^{1}(D^{\prime})$. Let $v_{\eta}$ be defined as
(20) $v_{\eta}(z)=(\xi,H_{D^{\prime}\setminus\eta}(z,\cdot))_{\mathcal{B}_{r}^{\prime}},\quad z\in D^{\prime}\setminus\eta.$
On $\left\{\mathrm{dist}(\eta,\mathcal{B}^{\prime})>\varepsilon\right\}$, it holds that $(u_{\xi}^{\varepsilon})_{\eta}+v_{\eta}=u_{\xi}$, and $((u_{\xi}^{\varepsilon})_{\eta},u_{\xi})_{\nabla(D^{\prime})}=0$ as $(u_{\xi}^{\varepsilon})_{\eta}$ vanishes on $\partial D^{\prime}$ and $u_{\xi}$ is harmonic in $D^{\prime}$. We can readily write:
$((u_{\xi}^{\varepsilon})_{\eta},(u_{\xi}^{\varepsilon})_{\eta})_{\nabla(D^{\prime})}=(v_{\eta},v_{\eta})_{\nabla(D^{\prime})}-(u_{\xi},u_{\xi})_{\nabla(D^{\prime})}.$
By harmonicity of $v_{\eta}$ and $u_{\xi}$, we have that
(21) $(v_{\eta},v_{\eta})_{\nabla(D^{\prime})}-(u_{\xi},u_{\xi})_{\nabla(D^{\prime})}=\int_{\mathcal{B}_{r}^{\prime}}(v_{\eta}\cdot\partial_{\nu}v_{\eta}-u_{\xi}\cdot\partial_{\nu}u_{\xi}).$
Using now the explicit expressions for $v_{\eta}$ and $u_{\xi}$ we obtain Equation (19).
We conclude the proof of the proposition by taking $\varepsilon\to 0$, as the terms computed above do not depend on $\varepsilon$. ∎
Remark 2.2.
A consequence of the proof is that $\mathrm{SLE}_{4}$ cannot be seen as a level-line of a single GFF, whose boundary values on the inner boundary components would not be functions of the level-line itself. Whereas we will not provide a proof, here is an intuitive argument: if a coupling $(\Phi_{D},\eta)$ of the sort were to exist, assuming that $D=\mathbb{A}_{p}$ for simplicity, we could condition on $\mathcal{A}_{+\lambda}$, the event that $\eta$ passes to the left of $C_{e^{-p}}$, and consider $D^{\prime}\subset\mathbb{A}_{p}$ a test domain passing to the left of $C_{e^{-p}}$ with $\mathrm{dist}(D^{\prime},C_{e^{-p}})>\varepsilon$. On $\mathbb{A}_{p-\delta}$, $\delta=\log\varepsilon$, we can write $\Phi_{\mathbb{A}_{p}}=\Phi_{\mathbb{A}_{p-\delta}}+u_{\xi}$ where $u_{\xi}$ is the harmonic extension of $\xi=\Phi_{\mathbb{A}_{p}}|_{C_{e^{-p}+\varepsilon}}$. The restriction property would then require that the level-line of $\Phi_{\mathbb{A}_{p-\delta}}$ conditioned to stay in $D^{\prime}$ is absolutely continuous w.r.t. the level-line of $\Phi_{D^{\prime}}^{g_{D^{\prime}}}$ with the right Radon-Nikodym derivative, and the proof shows that this is the case provided that the expectation of Equation (14) w.r.t. $\xi$ conditional on $\eta$ corresponds to $\exp(-m_{\mathbb{A}_{p}}(\eta,D^{\prime})/2)$. It is reasonable to imagine that this is only possible for all $D^{\prime}$ as above if $\xi\sim\mathcal{N}(g_{D,+\lambda},G_{D})$, so that taking $\varepsilon\to 0$ we recover that $\Phi_{D}|_{\mathcal{A}_{+\lambda}}=\Phi_{D}^{g_{D,+\lambda}}|_{\mathcal{A}_{+\lambda}}$.
The construction above determines the form of the partition function for non-crossing $\mathrm{SLE}_{4}$.
Corollary 12.
The partition function of $\mathrm{SLE}_{4}$ from $x$ to $y$ in $D$ can be expressed as
$Z_{D}(x,y):=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2})\cdot\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}}),$
and it is finite.
Let us explain and justify this expression. From the discussion in Remark 2.1 and the proof of Theorem 19, we may equivalently write $Z_{D}(x,y)$ as
$Z_{D}(x,y)$ $=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\exp(\frac{1}{2}(\norm{g_{D^{\prime}}}_{\nabla(D^{\prime})}^{2}-\norm{g_{\mathbf{b}}}_{\nabla(D)}^{2}))H_{D^{\prime}}(x,y)^{1/4}\cdot\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})$
$=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\exp(2\lambda^{2}(M(\mathcal{B}^{\prime}_{-},\mathcal{B}^{\prime}_{+})-M(\mathcal{B}_{\mathbf{b}}^{-},\mathcal{B}_{\mathbf{b}}^{+})))H_{D^{\prime}}(x,y)^{1/4}\cdot\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})$
for any test domain $D^{\prime}\subset D$, without needing to regularize the Dirichlet energies. The advantage of this expression is that conformal covariance such as in Equation (7) can be read directly, given that the probability term and the one corresponding to the difference of conformal moduli are conformally invariant, as mentioned in Section 1.3.
Remark 2.3.
In [IK13, HBB10], $\mathrm{SLE}_{4}$ is studied in the annulus. For a GFF in $(\mathbb{A}_{p},e^{ix},e^{iy})$, where $x,y\in\mathbb{R}$, with boundary conditions $-\lambda$ on $(e^{iy}e^{ix})$, $+\lambda$ on $(e^{ix}e^{iy})$ and $\mu$ on $C_{e^{-p}}\subset\partial\mathbb{A}_{p}$, it is shown that the corresponding level line is driven according to the annulus Loewner equation (see [Zha04]) with a driving process $\exp(iW_{t})$, where the relative position $g_{t}(e^{ix})\exp(-iW_{t})=:\exp(iB_{t})$ is such that $B_{t}$ is a Brownian bridge from (the representative in $[0,2\pi)$ modulo $2\pi$ of) $y-x$ to $\pi(1-\mu/\lambda)$, of time-duration $p$, and up to its exit time $T$ of $[0,2\pi]$. The event that $T=+\infty$ corresponds to the resulting level-line touching the inner boundary component $C_{e^{-p}}$ (recall that the conformal modulus of the slit domain plays the role of the time parametrization), whereas $T\leq p$ corresponds to the level-line reaching $e^{iy}$ from the left or right of $C_{e^{-p}}$, when $B_{T}=0$ or $2\pi$ respectively. Observe that we have $T\leq p$ almost surely when |$\absolutevalue{\mu}\geq\lambda$, as mentioned in Section 2.1.2.
In our case, the probabilities $\mathbb{P}_{\mathbb{A}_{p}}^{g_{-\lambda}}(\mathcal{A}_{-\lambda})$ and $\mathbb{P}_{\mathbb{A}_{p}}^{g_{+\lambda}}(\mathcal{A}_{+\lambda})$ correspond to the probability that a Brownian bridge from $y-x$ to $2\pi$ (resp. to 00) exits $[0,2\pi]$ through 00 (resp. $2\pi$), and the conformal modulus of the component of $\mathbb{A}_{p}\setminus\eta_{T}$ that is doubly connected is equal to $T$.
2.3. Crossing $\mathrm{SLE}_{4}$
We now go over the construction of the $\mathrm{SLE}_{4}$ measure in the crossing case, i.e. corresponding to a measure on simple curves joining two points belonging to two different connected components of the boundary of a multiply connected domain. There are two notable differences with the non-crossing case:
  • The relevant level-lines are coupled with multivalued GFFs, corresponding to the sum of a (Dirichlet, zero-boundary) GFF and a deterministic multivalued harmonic function with locally constant boundary conditions with a change of ±$\pm 2\lambda$ around $x$ and $y$ on $\partial D$, and a change of branch describing the winding of the curve along the different boundary components. In general these level-lines are again conditioned on topological events, but in the case of doubly connected domains the coupling can be realized directly with a single compactified GFF–a sum of a (Dirichlet, zero-boundary) GFF and multivalued harmonic function with random monodromy.
  • The topological events $\mathcal{A}_{\mathbf{b}},\mathbf{b}=(b_{1},\cdots,b_{n})\in\left\{\pm\lambda\right\}^{n}$, that amounted to assignations of $-\lambda$ or $+\lambda$ as boundary conditions of a GFF to the connected components of the boundary of the domain, correspond in the crossing case to events $\mathcal{A}_{\mathbf{b}},\mathbf{b}=(b_{\text{in}},b_{1},\dots,b_{n})\in\Lambda^{n+1}:=\mathbb{Z}\times(2\lambda\mathbb{Z}+\lambda)^{n}$. They represent the change in branch of the multivalued harmonic function acting as the mean of the GFF when approaching a boundary component (target or not). The component $b_{\text{in}}$ describes the winding of any $\eta\in\mathcal{A}_{\mathbf{b}}$ around $\mathcal{B}_{\text{in}}$.
[…]
We then let $\mathcal{A}_{\mathbf{b}}$ denote the set of simple curves $\eta:x\to y$ such that the parts of $\partial U$ that are contained between any two adjacent lifts of $\eta$ correspond to constant boundary values of $g_{\mathbf{b}}$. We finally define $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ as the law of the generalized level line of a GFF $\Phi_{D}$ in $D$ with boundary conditions given by $g_{\mathbf{b}}$ (see also Appendix A).
Figure 4. A configuration with a curve in $\mathcal{A}_{0,+3\lambda,+\lambda}$.
The point of the definition is just to guarantee that $g_{\mathbf{b}}$ corresponds to the boundary conditions $\lambda(1+2k)$ on every part of the boundary of $U$ between the lifts of $\eta$ starting from $2k\pi$ and $2(k+1)\pi$ (for any $\eta\in\mathcal{A}_{\mathbf{b}}$: see Figure 4).
The Dirichlet energy $\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}$ is then defined as $\norm{g_{\mathbf{b}}}_{\nabla(D^{\prime}),\text{reg}}$, where $D^{\prime}\subset U$ is any fundamental domain for the action of the translation by $2\pi$ in $U$. For instance, one may take $D^{\prime}$ as a parallelogram with vertices $-\pi,\pi,w-\pi,w+\pi$, for any $w$ such that $e^{iw}=y$, and regularize the divergences at 00 and $w$ like in Section 1.3.
Theorem 13.
Let $D\subset\mathbb{D}$ be a crossing admissible domain. The measure $\nu_{D}$ on simple curves from $x$ to $y$ in $D$ defined as
(22) $\nu_{D}(\eta):=\sum_{\mathbf{b}\in\Lambda^{n+1}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\mathbb{P}_{D}^{g_{\mathbf{b}}}(\eta)\cdot Z_{D,\mathbf{b}},$
where
$Z_{D,\mathbf{b}}=\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2}),$
is the measure of $\mathrm{SLE}_{4}$ in $(D,x,y)$.
Remark 2.4.
Here again, the expression for the partition function comes from
$\frac{Z_{D,\mathbf{b}}}{Z_{D^{\prime}}}=\exp(-\frac{1}{2}\left(\norm{g_{\mathbf{b}}}_{\nabla{(D^{\prime})}}^{2}-\norm{g_{D^{\prime}}}_{\nabla{(D^{\prime})}}^{2}\right)),$
that will appear in the proof, well-defined without regularization. This also justifies conformal invariance as in the non-crossing case.
Secondly, while it is necessary to construct the measures on the events $\mathcal{A}_{\mathbf{b}}$, the restriction for these events that the winding of a curve $\eta$ around $\mathcal{B}_{\text{in}}$ is given by $b_{\text{in}}$ is actually superfluous: indeed, the level line of $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ already almost surely satisfies this property, i.e. the monodromy of $g_{\mathbf{b}}$ fixes the winding of the level-line. This was proven in [IK13] for the case of the annulus, and extends to the general case by absolute continuity of the measures: see Appendix A.
In particular, when $D$ is a conformal annulus (corresponding to $n=0$), we get the following corollary:
Corollary 14.
When $D$ is a conformal annulus, $\nu_{D}$ corresponds to the law of a level-line of a compactified GFF $\Phi=\Phi_{D}^{0}+g$, where $g$ is a random multivalued harmonic function in $D$, independent of $\Phi_{D}^{0}$, taking values $g=g_{\mathbf{b}}$ with probability $Z_{D,\mathbf{b}}/\sum_{\mathbf{b}^{\prime}\in\mathbb{Z}}Z_{D,\mathbf{b}^{\prime}}$.
These probabilities can be computed explicitly, as we will do in a second Corollary: but first, we deduce from the Theorem an expression for the partition function for crossing $\mathrm{SLE}_{\kappa}$.
Corollary 15.
The partition function of crossing $\mathrm{SLE}_{\kappa}$ can be expressed as
$Z_{D}(x,y):=\sum_{\mathbf{b}\in\Lambda^{n+1}}\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2})$
and it is finite. When $D$ is a conformal annulus, we have
$Z_{D}(x,y)=\sum_{b_{\text{in}}\in\mathbb{Z}}\exp(-\frac{1}{2}\norm{g_{b_{\text{in}}}}_{\nabla(D),\text{reg}}^{2}).$
The summability of the series is provided by the exponential decay of $Z_{D,\mathbf{b}}$ as $\norm{\mathbf{b}}_{\infty}\xrightarrow[]{}\infty$. Indeed, one can write
$g_{\mathbf{b}}=g_{\left\{0,\lambda,\cdots,\lambda\right\}}+2\lambda b_{\text{in}}\cdot v_{b_{\text{in}}}+\sum_{i=1}^{n}b_{i}v_{i},$
[…]
We now turn back to computing the law of the winding of the curve under $\nu_{D}$.
Corollary 16.
If $\mathbb{P}$ denotes $\nu_{\mathbb{A}_{p}}^{\#}(x,y)$ and $\alpha$ is the representative of $(\arg y-\arg x)/(2\pi)$ in $[0,1)$, we have that
(23) $\mathbb{P}\left[\eta\in\mathcal{A}_{k}\right]=\mathbb{P}\left[\mathrm{Ind}_{\eta}(y)=k+\alpha\right]=\frac{Z_{D,k}}{Z_{D}}\varpropto e^{-\frac{\pi}{4}M(C_{e^{-p}},C_{1})\cdot(k+\alpha)^{2}}=e^{-\frac{\pi^{2}}{2p}\cdot(k+\alpha)^{2}},\quad k\in\mathbb{Z}.$
Proof.
As explained in Remark 2.4, the event $\mathcal{A}_{b_{\text{in}}}$ exactly corresponds to the event that $\mathrm{Ind}_{\eta}(y)=b_{\text{in}}+\alpha$ (with the winding defined consistently with $\alpha$), so that it suffices to compute $Z_{D,b_{\text{in}}}$ to identify the desired probabilities. Notice that $g_{b_{\text{in}}}=g_{0}+b_{\text{in}}f$, for $f=2\lambda\log\absolutevalue{z}/p$. Thus,
$\norm{g_{b_{\text{in}}}}_{\nabla(D),\text{reg}}^{2}$ $=\norm{g_{0}}_{\nabla(D),\text{reg}}^{2}+b_{\text{in}}^{2}\norm{f}_{\nabla(D)}^{2}+2b_{\text{in}}(g_{0},f)_{\nabla(D)}$
$=\norm{g_{0}}_{\nabla(D),\text{reg}}^{2}+b_{\text{in}}^{2}\cdot\frac{\pi^{2}}{p}+2b_{\text{in}}(g_{0},f)_{\nabla(D)}$
The term $(g_{0},f)_{\nabla(D)}$ is a function of $\alpha$, and the dependence is affine, given that $g_{0}$ can be constructed from the case $\alpha=0$ by precomposing by a rotation that is a function of the modulus of the input. Since at $\alpha=1$ it must be equal to $\pi^{2}/p$, we deduce that
$\norm{g_{b_{\text{in}}}}_{\nabla(D),\text{reg}}^{2}=\frac{\pi^{2}}{p}\cdot(b_{\text{in}}+\alpha)^{2}+C.$
This is also $4\lambda^{2}$ times the Dirichlet energy of the harmonic function equal to $1$ on $\mathcal{B}_{\text{in}}$ and 00 on $\mathcal{B}_{\text{out}}$, and the connection with the conformal modulus is provided by Section 1.3. ∎
Note that the asymptotics for Equation (23) are consistent with taking the size of $\mathcal{B}_{\text{in}}$ to zero (or $p\to\infty$ in $\mathbb{A}_{p}$) as in this case the winding is known to be infinite almost surely [Sch00].
Proof of Theorem 13.
Let $\tilde{\nu}_{D}$ be the measure in the statement of the Theorem, $D^{\prime\prime}\subset D$ be a (crossing) test domain, and let $\mathbf{b}^{\prime}\in\Lambda^{n+1}$ be such that $\left\{\eta\in D^{\prime\prime}\right\}\subset\mathcal{A}_{\mathbf{b}^{\prime}}$. This time again, instead of comparing $\tilde{\nu}_{D}$ directly to $\nu_{D^{\prime\prime}}(x,y)$, we use an intermediary domain $D^{\prime}=D\setminus\delta$, where $\delta$ is the arc of $\partial D^{\prime\prime}$ from the right of $x$ to $\mathcal{B}_{\text{in}}$. In particular, $\mathrm{SLE}_{4}$ from $x$ to $y$ in $D^{\prime}$ is the non-crossing $\mathrm{SLE}$ that was previously constructed ($D^{\prime}$ is conformally equivalent to a non-crossing admissible domain), and its law $\nu_{D^{\prime}}(x,y)$ is described by Theorem 9, in which case it holds that
$\frac{\mathrm{d}\nu_{D^{\prime\prime}}(x,y)}{\mathrm{d}\nu_{D^{\prime}}(x,y)}(\eta)=\mathbf{1}_{\eta\subset D^{\prime\prime}}\exp(\frac{1}{2}m_{D^{\prime}}(\eta,D^{\prime}\setminus D^{\prime\prime})).$
The set of $\mathbf{b}\in\Lambda^{n+1}$ such that $\left\{\eta\in D^{\prime}\right\}\cap\mathcal{A}_{\mathbf{b}}\neq\emptyset$ can be naturally identified with {±$\left\{\pm\lambda\right\}^{n}$, as ${\eta\in D^{\prime}}$ fixes the winding around $\mathcal{B}_{\text{in}}$, and leaves the freedom of going to the left or the right of each $\mathcal{B}_{i}$ in $D^{\prime}$. Let us under this identification consider $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$, and denote $g_{D^{\prime}}=g_{D^{\prime},\mathbf{b}},g_{D}=g_{D,\left\{b_{\text{in}},\mathbf{b}\right\}}$ (where $b_{\text{in}}$ is determined by $D^{\prime}$) for simpler notation. We have to prove that $\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}$ is absolutely continuous with respect to $\mathbb{P}_{D}^{g_{D}}$ with Radon-Nikodym derivative given by
$\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}}{\mathrm{d}\mathbb{P}_{D}^{g_{D}}}(\eta)=\mathbf{1}_{\eta\subset D^{\prime}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\frac{Z_{D,\mathbf{b}}}{Z_{D^{\prime},\mathbf{b}}}\exp(\frac{1}{2}m_{D}(\eta,D\setminus D^{\prime})).$
This is achieved once again by computing the average of the Radon-Nikodym derivative for fixed $\xi$ over $\xi\sim\mathcal{N}(0,G_{D\setminus\eta}|_{\delta})$. Notice that conditionally on $\eta$, $\Phi_{D}^{g_{D}}|_{D^{\prime}}$ can be written $\Phi_{D\setminus\eta}^{0}+u_{\xi}+h_{\eta}$ where $u_{\xi}$ is the harmonic extension of $\xi$ to $D^{\prime}$, $h_{\eta}$ is multivalued with every branch in $D^{\prime}$ of the form $(2n-1)\lambda+2\lambda\mathbf{1}_{\text{right of }\eta}(z)$, so that the law $\mathbb{P}_{D}^{g_{D}}$ knowing $\xi$ and restricted to the event $\mathcal{A}_{\mathbf{b}}\cap\left\{\eta\subset D^{\prime}\right\}$ is the same as $\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}},\xi}$ from the non-crossing case thanks to the construction of $g_{D}$ and the definition of $\mathcal{A}_{\mathbf{b}}$. Similarly to the non-crossing case, we start with the computation of the Radon-Nikodym derivative for $\xi$ fixed. Although $\delta$ is one single crosscut, it is part of the boundary of both $D_{l}^{\prime}$ and $D_{r}^{\prime}$, the left and right connected components of $D^{\prime}$ relative to $\eta$: as such, we denote formally by $\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}^{\prime}$) $\delta$ seen as the boundary of $D_{l}^{\prime}$ (resp. $D_{r}^{\prime}$). We will be brought to consider Brownian excursions from $\delta$ to $\delta$ in $D$, which we understand as excursions between the two preimages of $\delta$ by $x\mapsto e^{ix}$, that we identify to $\mathcal{B}_{l}^{\prime}$ and $\mathcal{B}_{r}^{\prime}$. Denote also by $\nu_{l}$ (resp. $\nu_{r}$) the normal vector to $D_{l}^{\prime}$ (resp $D_{r}^{\prime}$) at $x\in\delta$.
Proposition 17.
For all $\xi\in H^{-1}(\delta)$, it holds that $\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}\ll\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}},\xi}$, and we have explicitly that
(24) $\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}}{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}},\xi}}(\eta)\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\in D^{\prime}}=\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\in D^{\prime}}\exp(\frac{1}{2}(\xi,M\xi)_{\delta}-((\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}},\xi)_{\delta}),$
where
$M=(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}}+(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}+2(H_{\partial D^{\prime}})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{r}^{\prime}}$
with the subscripts/superscripts $\mathcal{B}_{l}^{\prime},\mathcal{B}_{r}^{\prime}$ to differentiate when the operators correspond to Brownian excursions from $\mathcal{B}_{i}^{\prime}$ to $\mathcal{B}_{j}^{\prime}$, $i,j=l,r$.
We first finish the proof of the theorem before going over the proof of the proposition. When computing the average of Equation (24) with Lemma 6, the determinant term is that of
(25) $I-G_{D\setminus\eta}|_{\delta}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}}-G_{D\setminus\eta}|_{\delta}(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}-2G_{D\setminus\eta}|_{\delta}(H_{\partial D^{\prime}})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{r}^{\prime}}.$
This can again be understood with another decomposition formula for Brownian loops. One can further decompose the expressions obtained after applying Lemma 1 by distinguishing on which side the Brownian path leaves and hits $\delta$:
$G_{D}-G_{D\setminus\eta}$ $=(H_{D\setminus\eta})_{\eta}^{\delta}(H_{D^{\prime}})_{\delta}^{\eta}G_{D}$
$=G_{D\setminus\eta}\left((H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{l}^{\prime}}-(H_{\partial D^{\prime}}-H_{\partial(D^{\prime}\setminus\eta)})_{\mathcal{B}_{r}^{\prime}}^{\mathcal{B}_{r}^{\prime}}-2(H_{\partial D^{\prime}})_{\mathcal{B}_{l}^{\prime}}^{\mathcal{B}_{r}^{\prime}}\right)G_{D}$
From here, Proposition 2 once again provides that the determinant term in the Gaussian integration is equal to
$m_{D}(\eta,D\setminus D^{\prime}).$
This also shows that
$G_{D\setminus\eta}^{1/2}(I-G_{D\setminus\eta}^{1/2}MG_{D\setminus\eta}^{1/2})^{-1}G_{D\setminus\eta}^{1/2}=G_{D}$
and the exponential term is equal to
(26) $\exp(\frac{1}{2}((\partial_{\nu_{l}}+\partial_{\nu_{r}})(g_{D^{\prime}}),G_{D}((\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}}))_{\delta})=\exp(-((\partial_{\nu_{l}}+\partial_{\nu_{r}})(g_{D^{\prime}}-g_{D}),G_{D}((\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}}))_{\delta}),$
given that the gradient of $g_{D}$ is single valued. Now we need to invert $G_{D}|_{\delta}$, i.e. the operator
$f\mapsto\int_{\delta}G_{D}(\cdot,x)f(x)\mathrm{d}x$
that from a function in $L^{2}(\delta)$ outputs a continuous function in $D$, harmonic in $D\setminus\delta$, with values 00 on $\partial D$ and whose normal derivative jumps by $f(x)$ across $\delta$ at $x$. If one now replaces $g_{D^{\prime}}$ with the multivalued function $g_{D^{\prime},m}$ that is equal to $g_{D}+2n\lambda$ in every branch in $D^{\prime}$, then $g_{D}-g_{D^{\prime},m}$ is single-valued and $g_{D}-g_{D^{\prime},m}$ satisfies the conditions above [^3] In fact the inverse of the Green’s function here is also a Neumann-type operator that computes the difference of the normal derivatives of the harmonic extensions of $f$ on both sides of $\delta$.: it follows that $G_{D}((\partial_{\nu_{l}}+\partial_{\nu_{r}})(g_{D}-g_{D^{\prime},m}))=(g_{D}-g_{D^{\prime},m}))|_{\delta}$. We rewrite Equation (26) as
$\exp(\frac{1}{2}((g_{D}-g_{D^{\prime},m}),(\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}})_{\delta})=\exp(\frac{1}{2}\int_{D^{\prime}}g_{D}\cdot\nabla g_{D^{\prime}}-\frac{1}{2}\norm{\nabla g_{D^{\prime}}}_{D^{\prime}}^{2})$
and to conclude it suffices to note that
$\int_{D^{\prime}}g_{D}\cdot\nabla g_{D^{\prime}}=\int_{\delta}g_{D^{\prime}}\cdot(\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D}=\int_{\delta}g_{D}\cdot(\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D}=\norm{\nabla g_{D}}_{D^{\prime}}^{2}$
given that $g_{D}$ has monodromy $2\lambda$. ∎
We now turn to the proof of Proposition 17.
Proof of Proposition 17.
The proof starts identically to that of Proposition 11, up to the computation of $((u_{\xi}^{\varepsilon})_{\eta},(u_{\xi}^{\varepsilon})_{\eta})_{\nabla(D^{\prime})}$. Write
$u_{\xi}(z)=\int_{\delta}H_{D\setminus\delta}(z,\mathrm{d}y)\xi(y)=\underbrace{\int_{\mathcal{B}_{l}^{\prime}}H_{D^{\prime}}(z,\mathrm{d}y)\xi(y)}_{=:u_{l}(z)}+\underbrace{\int_{\mathcal{B}_{r}^{\prime}}H_{D^{\prime}}(z,\mathrm{d}y)\xi(y)}_{{=:u_{r}(z)}}$
We still have that $u_{\xi}=(u_{\xi}^{\varepsilon})_{\eta}+v_{\eta}$, with $v_{\eta}$ defined as
$v_{\eta}(z)=\int_{\mathcal{B}_{l}^{\prime}}H_{D_{l}^{\prime}}(z,\mathrm{d}y)\xi(y)+\int_{\mathcal{B}_{r}^{\prime}}H_{D_{r}^{\prime}}(z,\mathrm{d}y)\xi(y).$
Similarly to the proof of Proposition 11, assume that $\xi$ is smooth, and compute
$((u_{\xi}^{\varepsilon})_{\eta},(u_{\xi}^{\varepsilon})_{\eta})_{\nabla(D^{\prime})}$ $=(v_{\eta},v_{\eta})_{\nabla(D^{\prime})}-(u_{l},u_{l})_{\nabla(D^{\prime})}-(u_{r},u_{r})_{\nabla(D^{\prime})}+2(u_{l},\partial_{\nu}u_{r})_{\mathcal{B}^{\prime}}$
$=\int_{\mathcal{B}_{l}^{\prime}}\xi\cdot(H_{\partial D^{\prime}}-H_{\partial D_{l}^{\prime}})^{\mathcal{B}_{l}^{\prime}}\xi+\int_{\mathcal{B}_{r}^{\prime}}\xi\cdot(H_{\partial D^{\prime}}-H_{\partial D_{r}^{\prime}})^{\mathcal{B}_{r}^{\prime}}\xi+2\int_{\mathcal{B}_{r}^{\prime}}\xi\cdot(H_{\partial D^{\prime}})^{\mathcal{B}_{r}^{\prime}}\xi,$
seeing $\xi$ as an element of $H^{-1}(\mathcal{B}_{l}^{\prime})$ or $H^{-1}(\mathcal{B}_{r}^{\prime})$.
The computation of the term $(\Phi_{\eta},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}$ is identical to Equation (18), taking into account that $\mathcal{B}_{l}^{\prime}$ and $\mathcal{B}_{r}^{\prime}$ are seen as distinct parts of the boundary of $D^{\prime}$:
$(\Phi_{\eta},u_{\xi}^{\varepsilon})_{\nabla(D^{\prime})}=-\int_{\partial D^{\prime}}\xi\cdot\partial_{\nu}h_{\eta}=-\int_{\mathcal{B}_{l}^{\prime}}\xi\cdot\partial_{\nu}h_{\eta}-\int_{\mathcal{B}_{r}^{\prime}}\xi\cdot\partial_{\nu}h_{\eta}=(\partial_{\nu}g_{D^{\prime}},\xi)_{\mathcal{B}_{l}^{\prime}}+(\partial_{\nu}g_{D^{\prime}},\xi)_{\mathcal{B}_{r}^{\prime}}.$
3. Construction of $\mathrm{SLE}_{\kappa}$ in multiply connected domains
We now turn to a more general construction of $\mathrm{SLE}_{\kappa}$ in multiply connected domains for the whole range $\kappa\in(8/3,4]$, arguably simpler than the previous construction of $\mathrm{SLE}_{4}$, but specific to the non-crossing case. We once again start by recalling classical facts about restriction measures and CLEs.
3.1. Restriction measures and CLEs
Consider $D\subset\mathbb{C}$ a (non-crossing) admissible domain and a Poisson point process (PPP) of intensity $h$ with respect to the (infinite) measure of Brownian excursions from $\mathcal{B}_{l}$ in $D$
$\pi\iint_{\mathcal{B}_{l}\times\mathcal{B}_{l}}\mu_{\partial D}(z,w)\mathrm{d}z\mathrm{d}w.$
Denote by $R_{h}^{D,\mathcal{B}_{l}}$ its filling, i.e. the complement of the connected component of $D$ minus a sample of the PPP containing $\mathcal{B}_{r}$. Note that the factor of $\pi$ is due to our convention for the Poisson kernel.
Lemma 18.
In the setting above and for $D^{\prime}\subset D$ such that $D\setminus D^{\prime}$ is at positive distance from $\mathcal{B}_{l}$, the probability that all loops in $R_{h}^{D,\mathcal{B}_{l}}$ stay in $D^{\prime}$ is equal to
(27) $\exp(\frac{h\pi}{4}(\norm{u_{D^{\prime}}}_{\nabla(D^{\prime})}^{2}-\norm{u_{D}}_{\nabla(D)}^{2})),$
where $u_{D}$ (resp. $u_{D^{\prime}}$) is harmonic with boundary values $-1$ on $\mathcal{B}_{l}$, $+1$ on the rest of $\partial D$ (resp. $\partial D^{\prime}$).
Proof.
By definition of the PPP and the observation that the excursion measures satisfy a form of restriction (see Section 1.2), the random variable that counts the loops in a realization of the PPP that exit $D^{\prime}$ is Poisson with parameter
$h\pi\iint_{\mathcal{B}_{l}^{\prime}}(H_{\partial D}(x,y)-H_{\partial D^{\prime}}(x,y))\mathrm{d}x\mathrm{d}y.$
As in Section 1.3, this might be rewritten as
$\frac{h\pi}{4}(\norm{u_{D^{\prime}}}_{\nabla(D^{\prime})}^{2}-\norm{u_{D}}_{\nabla(D)}^{2}).\qed$
Naturally the same construction can be replicated ad hoc replacing $\mathcal{B}_{l}$ by $\mathcal{B}_{\mathbf{b}}^{-}$ for $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$. Importantly, as mentioned in Remark 2.1 it can be verified that the regularized Dirichlet energy satisfies the following conformal covariance property, for $\varphi:D\to\varphi(D)$ a conformal equivalence:
$\exp(-\frac{1}{2}\norm{\varphi(u_{D})}_{\nabla(D),\text{reg}}^{2})=\absolutevalue{\varphi^{\prime}(x)}^{2/\pi}\absolutevalue{\varphi^{\prime}(y)}^{2/\pi}\exp(-\frac{1}{2}\norm{u_{\varphi(D)}}_{\nabla(\varphi(D)),\text{reg}}^{2}),$
so that up to a multiplicative constant,
$\exp(\norm{u_{D}}_{\nabla(D),\text{reg}}^{2})=H_{\partial D}(x,y)^{-4/\pi}$
when $D$ is simply connected. This is particularly important in the case above, as Equation (27) becomes
$\left(\frac{H_{\partial D^{\prime}}(x,y)}{H_{\partial D}(x,y)}\right)^{h},$
i.e. $R_{h}^{D,\mathcal{B}_{l}}$ is a one-sided restriction measure of exponent $h$ (see [LSW03] for more about restriction measures).
On the other hand, the Conformal Loop Ensembles (CLE) are another one-parameter family of conformally invariant measures, this time on collections of non-nested disjoint simple loops in a domain $D$. There are various ways to construct them (see [SW12]), but we will see them as the collection of the outermost boundaries of clusters in a Brownian loop-soup (i.e. a Poisson point process with respect to the Brownian loop measure) sample with intensity $\mathbf{c}/2=\mathbf{c}(\kappa)/2$, $0<\mathbf{c}\leq 1$. We stress that the intensity parameter in [SW12, Law09] is missing this factor of $1/2$: see also [Lup18] for a more thorough explanation. Note that $1/2$ is the critical identity, above which there is a.s. one cluster in the Brownian loop soup and its boundary coincides with $\partial D$, and that the range $8/3<\kappa\leq 4$ exactly corresponds to $0<\mathbf{c}\leq 1$.
The CLE also satisfies a form of conformal restriction: for $D^{\prime}\subset D$, consider the set $\Gamma^{\prime}$ obtained from a realization $\Gamma$ of the CLE in $D$ consisting of all the loops that intersect $D^{\prime}$, and let $D_{\Gamma}^{\prime}$ be the random set given by the union of $D^{\prime}$ with the filling of $\Gamma$. Then, conditionally on $D_{\Gamma}^{\prime}$, the law of the loops of the CLE that do not intersect $D_{\Gamma}^{\prime}$ is that of a CLE in $D\setminus D_{\Gamma}^{\prime}$.
3.2. The non-crossing $\mathrm{SLE}_{\kappa}$
We work with the notation of the previous section, only redefining the parameters $\mathbf{b}$ to belong to {±$\left\{\pm 1\right\}^{n}$ (the purpose of the factor $\lambda$ was just to make a notational connection with the boundary conditions of the GFFs, but all that matters is the sign of $\mathbf{b}$). We will also consider for $D$ an admissible domain and $\mathbf{b}\in\left\{\pm 1\right\}^{n}$ the harmonic function $u_{D}$ given by the boundary values $-1$ on $\mathcal{B}_{\mathbf{b}}^{-}$, $+1$ on $\mathcal{B}_{\mathbf{b}}^{+}$, so that $u_{\mathbf{b}}=g_{\mathbf{b}}/\lambda$ from the previous section.
Consider $K$ a sample of the restriction measure $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$, $\Gamma$ a sample of an independent $\operatorname{CLE}_{\kappa}^{D}$, and $C$ the union of $K$ with all the loops of $\Gamma$ that it intersects. We let $\mathbb{P}_{D}^{\mathbf{b}}$ be the law of the right boundary of the component of $C$ that contains the arc $(yx)$, i.e. the boundary of its filling minus $(yx)$. This is a simple curve in $D$ (see [WW13] for an argument in the simply connected case: the general case is identical). In other words, $\mathbb{P}_{D}^{\mathbf{b}}=F_{\ast}(\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}})$, where $F$ identifies the path according to the construction above given a realization of the $\operatorname{CLE}$ and of the restriction measure.
Theorem 19.
Let $D\subset\mathbb{D}$ be admissible and $8/3<\kappa\leq 4$. The measure $\nu_{D}$ on simple curves from $x$ to $y$ in $D$ defined as
(28) $\nu_{D}(\eta):=\sum_{\mathbf{b}\in\left\{\pm 1\right\}^{n}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\mathbb{P}_{D}^{\mathbf{b}}(\eta)\cdot Z_{D,\mathbf{b}},$
where
(29) $Z_{D,\mathbf{b}}=\exp(-\frac{\pi h}{4}\norm{u_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2}),$
is the measure of $\mathrm{SLE}_{\kappa}$ in $(D,x,y)$.
Note that $\pi h_{4}/4=\pi/16=\lambda^{2}/2$ so that the expression above is compatible with Equation (10).
From this construction we can infer the expression of the partition function for non-crossing $\mathrm{SLE}_{\kappa}$.
Corollary 20.
The partition function of $\mathrm{SLE}_{\kappa}$ from $x$ to $y$ in $D$, $8/3<\kappa\leq 4$, can be expressed as
$Z_{D}(x,y):=\sum_{\mathbf{b}\in\left\{\pm 1\right\}^{n}}\exp(-\frac{\pi h_{\kappa}}{4}\norm{u_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2})\cdot\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})$
and it is finite.
The proof of the Theorem will again follow Lawler’s characterization of the $\mathrm{SLE}$ measures. In fact, it will be sufficient to show that for $K\subset D$ a compact at positive distance from $\mathcal{B}_{l}$ such that $\mathcal{B}_{r}\cup K$,
(30) $\frac{\mathrm{d}\nu_{D\setminus K}(x,y)}{\mathrm{d}\nu_{D}(x,y)}(\eta)=\mathbf{1}_{\eta\subset D}\exp(\frac{\mathbf{c}}{2}m_{D}(\eta,K)).$
Indeed, this would characterize $\nu_{D}$ seen as a law on connected hulls containing $(yx)$, i.e. right-sided restriction measures from $x$ to $y$: since the $\mathrm{SLE}_{\kappa}$ measures should have the same Radon-Nikodym derivative, they must correspond to the same right-sided restriction measure, and the fact that their boundary are given by a simple curve is enough to assert that they also coincide as measures on simple curves from $x$ to $y$.
Proof.
Let $D^{\prime}:=D\setminus K$ be as above. Up to relabeling the boundary components $(\mathcal{B}_{i})_{1\leq i\leq n}$, we may assume that $(\mathcal{B}_{i})_{1\leq i\leq k}$ are the inner boundary components of $D^{\prime}$. We have to verify that $\nu_{D}$ satisfies Equation (30), i.e. that
(31) $\frac{\mathrm{d}\nu_{D^{\prime}}}{\mathrm{d}\nu_{D}}(\eta)=\sum_{\begin{subarray}{c}\mathbf{b}=(\mathbf{b}_{k},1,\dots,1)\\ \mathbf{b}_{k}\in\left\{\pm 1\right\}^{n-k}\end{subarray}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{\mathbf{b}_{k}}}{\mathrm{d}\mathbb{P}_{D}^{\mathbf{b}}}(\eta)\frac{Z_{D^{\prime},\mathbf{b}_{k}}}{Z_{D,\mathbf{b}}}=\sum_{\begin{subarray}{c}\mathbf{b}=(\mathbf{b}_{k},1,\dots,1)\\ \mathbf{b}_{k}\in\left\{\pm 1\right\}^{n-k}\end{subarray}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}\exp(\frac{\mathbf{c}}{2}m_{D}(\eta,D\setminus D^{\prime})).$
Because on $\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}$, $\mathbb{P}_{D}^{\mathbf{b}}$ (resp. $\mathbb{P}_{D^{\prime}}^{\mathbf{b}_{k}}$) is a deterministic function of a realization of $\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ (resp. of $\operatorname{CLE}_{\kappa}^{D^{\prime}}\otimes R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}$), it suffices to compute [^4] We are abusing notation: the equation should be understood as $\frac{\mathrm{d}\operatorname{CLE}_{\kappa}^{D^{\prime}}\otimes R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}(F(\Gamma,K))=\frac{\mathrm{d}\operatorname{CLE}_{\kappa}^{D^{\prime}}}{\mathrm{d}\operatorname{CLE}_{\kappa}^{D}}(F(\cdot,K))|_{\cdot=\Gamma}\frac{\mathrm{d}R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}(F(\Gamma,\cdot))|_{\cdot=K}$
$\frac{\mathrm{d}\operatorname{CLE}_{\kappa}^{D^{\prime}}\otimes R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}\operatorname{CLE}_{\kappa}^{D}\otimes R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}(\eta)=\frac{\mathrm{d}\operatorname{CLE}_{\kappa}^{D^{\prime}}}{\mathrm{d}\operatorname{CLE}_{\kappa}^{D}}(\eta)\frac{\mathrm{d}R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}(\eta).$
On the one hand, the sample of the Brownian loop soup in $D$ of intensity $\mathbf{c}/2$ from which $\operatorname{CLE}_{\kappa}^{D}$ is constructed consists of loops in $D^{\prime}$ and loops intersecting $D^{\prime}$. The former corresponds to a sample of a Brownian loop soup in $D^{\prime}$ and the outermost boundaries of its clusters yield a realization of $\operatorname{CLE}_{\kappa}^{D^{\prime}}$. One can therefore construct a curve $\tilde{\eta}^{\prime}$ from $\operatorname{CLE}_{\kappa}^{D^{\prime}}$, and then explore the boundary of the loops in the remainder of the loop soup in $D$ to obtain $\tilde{\eta}$. Both curves are then such that either $\tilde{\eta}^{\prime}=\tilde{\eta}$, or $\tilde{\eta}\not\subset D^{\prime}$: and the former happens if and only if the loop soup does not contain loops intersecting both $\eta$ and $D\setminus D^{\prime}$. Since the number of these loops is a Poisson random variable with parameter $\mathbf{c}\cdot m_{D}(\eta,D\setminus D^{\prime})/2$, we obtain that
$\frac{\mathrm{d}\operatorname{CLE}_{\kappa}^{D^{\prime}}}{\mathrm{d}\operatorname{CLE}_{\kappa}^{D}}(\eta)=\mathbb{P}\left[\not\exists\gamma\in\operatorname{CLE}_{\kappa}^{D}:\gamma\cap\eta\neq\emptyset,\gamma\cap(D\setminus D^{\prime})\neq\emptyset\right]^{-1}=\exp(\frac{\mathbf{c}}{2}m_{D}(\eta,D\setminus D^{\prime})).$
On the other hand, $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ can be obtained by first sampling $R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}$, then an independent Poisson point process of Brownian excursions intersecting $D\setminus D^{\prime}$, of intensity $h=h_{\kappa}$. Again here, the curve $\tilde{\eta}^{\prime}$ obtained from $R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}$ and the curve $\eta$ obtained from $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ are such that $\tilde{\eta}=\tilde{\eta}^{\prime}$ or $\tilde{\eta}\not\subset D^{\prime}$, the former happening if and only if the realisation of the PPP does not contain excursions intersecting $\mathcal{B}_{r}^{\prime}$ (note that these excursions necessarily hit $\eta$), i.e. if and only if no excursion in the realization of $R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}$ hit $D\setminus D^{\prime}$. Once again we find that
$\frac{\mathrm{d}R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}}{\mathrm{d}R_{h}^{D^{\prime},\mathcal{B}_{\mathbf{b}}^{-}}}(\eta)=\mathbb{P}\left[\not\exists\gamma\in R_{h}^{D,\mathcal{B}_{\mathbf{b}}^{-}}:\gamma\cap D\setminus D^{\prime}\neq\emptyset\right]^{-1}=\exp(\frac{\pi h}{4}(\norm{g_{D^{\prime},\mathbf{b}_{k}}}_{\nabla(D^{\prime})}^{2}-\norm{g_{D,\mathbf{b}}}_{\nabla(D)}^{2})),$
using Lemma 18. The right-hand term of this equality cancels out with $Z_{D^{\prime},\mathbf{b}_{k}}/Z_{D,\mathbf{b}}$ and concludes the proof. ∎
Appendix A Existence of level-lines for compactified GFF on crossing admissible domains
Here we explain in more details the existence of level-lines for compactified GFF, in the crossing case. Let $(D,x,y)$ be a crossing admissible domain: suppose first that $D$ is one-connected, and by conformal invariance consider $D=\mathbb{A}_{p}$. Let $\mathbf{b}=b_{\text{in}}\in\mathbb{Z}$ and $g_{b_{\text{in}}}$ as above, the multivalued harmonic function with locally constant boundary conditions with a jump of $+2\lambda$ across $x$ and $y$, and a change of fixed branch of $2\lambda b_{\text{in}}$ from the outer to the inner boundary of $D$. The existence of a level-line for a (compactified) GFF $\Phi_{\mathbb{A}_{p}}^{g_{b_{\text{in}}}}$ is then provided by the following result:
Proposition 21 ([IK13, Proposition 10]).
Let $\Phi_{\mathbb{A}_{p}}^{0}$ be a (zero Dirichlet) GFF in $\mathbb{A}_{p}$ and $\Phi_{\mathbb{A}_{p}}^{g_{b_{\text{in}}}}:=\Phi_{\mathbb{A}_{p}}^{0}+g_{b_{\text{in}}}$. Then there exists a local set coupling $(\Phi_{\mathbb{A}_{p}}^{g_{b_{\text{in}}}},\eta,h_{\eta})$ where $h_{\eta}$ is multivalued in $\mathbb{A}_{p}\setminus\eta$ with constant branches equal to $\lambda+2n\lambda$, and $\eta$ is an annulus $\mathrm{SLE}_{4}$ in $\mathbb{A}_{p}$, i.e. a Loewner chain for the annulus Loewner equation [Zha04] with driving function corresponding to a semimartingale with an explicit drift. The winding of $\eta$ at $y$ is almost surely equal to $b_{\text{in}}+\alpha$, with $\alpha$ the representative of $(\arg y-\arg x)/(2\pi)$ in $[0,1)$.
The proof is very similar to the case of the usual level line coupling in [SS13, Dub09] and relies on finding the right martingales to then use Itô’s formula: a sufficient condition for the law of the GFF to be equal conditionally on $\eta_{t}$ to a GFF in $\mathbb{A}_{p}\setminus\eta_{t}$ with mean given by $g_{b_{\text{in}},t}:=f_{t}^{-1}(g_{b_{\text{in}}})$, where $f_{t}:\mathbb{A}_{p}\to\mathbb{A}_{p-t}$ is the mapping-out function (recall that the modulus plays the role of the time-parametrization), is that for all $z\in\mathbb{A}_{p}$, $\left\{g_{t}(z)\right\}_{0\leq t\leq p}$ is a (continuous) uniformly bounded martingale with quadratic cross variation given by $G_{\mathbb{A}_{p}}(z,z^{\prime})-G_{\mathbb{A}_{p-\cdot}}(z,z^{\prime})$. In other words, $G_{\mathbb{A}_{p-\cdot}}(z,z^{\prime})$ should be a process of finite variation with $g_{\cdot}(z)g_{\cdot}(z^{\prime})-G_{\mathbb{A}_{p-\cdot}}(z,z^{\prime})$ a martingale. This guarantees that the characteristic function of the two GFFs mentioned above agree. Itô’s formula then provides differential equations that the drift of the driving function should satisfy in order for the conditions above to be met.
In fact, the form of the coupling that is established in [IK13, Proposition 10] is weaker than a local set coupling, but this is because their set-up (see their Theorem 2) encompasses general flow-lines where $h_{\eta}$ is possibly unbounded in a neighborhood of $\eta$, which is not the case above; in the case of level lines the local set coupling can be achieved as in [SS13].
When $D$ is $n$-connected with $n\geq 1$, existence of level-lines follows by absolute continuity: the argument is almost identical to the proof of Lemma 15 in [ASW19]. By conformal invariance we can take $D$ as an annulus $\mathbb{A}_{p}$ minus a set $K$ given by a finite union of closed disks. Let also $g_{D}$ be a multivalued harmonic function in $D$ with locally constant boundary conditions across $x$ and $y$, a change of fixed branch of $2\lambda b_{\text{in}}$ from the outer to the inner boundary of $D$, arbitrary values on $\partial K$, and such that $g(re^{i\theta+i2\pi})=g(re^{i\theta})+2\lambda$. Given $D^{\prime}\subset D$ a test domain, one can define the level-line $\eta$ of $\Phi_{D}^{g_{D}}:=\Phi_{D}^{0}+g_{D}$ up until its exit time of $D^{\prime}$ as the level-line in $D^{\prime}$ of a GFF $\Phi_{D^{\prime}}^{0}+g_{D}|_{D^{\prime}}$. Uniqueness of this level-line implies that they are compatible so that $\eta$ is well-defined (and actually a measurable function of the field). Similarly, if all branches of $g_{D}$ are such that |$\absolutevalue{g_{D}}\geq\lambda$ on $\partial K$, then almost surely $\eta$ doesn’t hit $\partial K$: the argument is here almost identical to Lemma 16 in [ASW19]. It therefore almost surely ends at $y$ with a winding described by $b_{\text{in}}$ since for all $D^{\prime}\subset D$ a test domain, the law of $\eta$ up to its exit time of $D^{\prime}$ is absolutely continuous with respect to that of the level-line $\tilde{\eta}$ of $\Phi_{\mathbb{A}_{p}}^{g_{b_{\text{in}}}}$, which ends at $y$ with winding described by $b_{\text{in}}$ on the event of positive probability that it does not exit $D^{\prime}$.
Appendix B Proof of Lemma 10
Proof of Lemma 10.
Let $f\in C_{b}(E)$. Denoting $R_{Y_{2},y}(x)=\mathrm{d}\mathbb{P}_{X}/\mathrm{d}\mathbb{P}_{Y_{1}|Y_{2}=y}(x)$, we compute:
$\int_{E}f(x)\mathrm{d}\mathbb{P}_{X}(x)=\int_{E}\int_{E}f(x)\mathrm{d}\mathbb{P}_{X}(x)\mathrm{d}\mathbb{P}_{Y_{2}}(y)=\int_{E}\int_{E}f(x)R_{Y_{2},y}(x)\mathrm{d}\mathbb{P}_{Y_{1}|Y_{2}=y}(x)\mathrm{d}\mathbb{P}_{Y_{2}}(y).$
Using Bayes’ formula, we rewrite the last term as
$\int_{E}\int_{E}f(x)R_{Y_{2},y}(x)\mathrm{d}\mathbb{P}_{Y_{1}|Y_{2}=y}(x)\mathrm{d}\mathbb{P}_{Y_{2}}(y)=\int_{E}\int_{E}f(x)R_{Y_{2},y}(x)\mathrm{d}\mathbb{P}_{Y_{2}|Y_{1}=x}(y)\mathrm{d}\mathbb{P}_{Y_{1}}(x).$
But on the other hand, denoting $R(x)=\mathrm{d}\mathbb{P}_{X}/\mathbb{P}_{Y_{1}}(x)$, we have that
$\int_{E}f(x)\mathrm{d}\mathbb{P}_{X}(x)=\int_{E}f(x)R(x)\mathrm{d}\mathbb{P}_{Y_{1}}(x),$
from which we deduce that
$R(x)=\int_{E}R_{Y_{2},y}(x)\mathrm{d}\mathbb{P}_{Y_{2}|Y_{1}=x}(y).$